### v2

# Steiner Traveling Salesman Problem with Time Windows and Pickup–Delivery: integrating classical and quantum optimization

Abstract
We propose the Steiner Traveling Salesman Problem with Time Windows and Pickup and Delivery, an advanced and practical extension of classical routing models. This variant integrates the characteristics of the Steiner Traveling Salesman Problem with time-window constraints, pickup and delivery operations and vehicle capacity limitations. These features closely mirror the complexities of contemporary logistics challenges, including last-mile distribution, reverse logistics and on-demand service scenarios. To tackle the inherent computational difficulties of this NP-hard problem, we propose two specialized mathematical formulations: an arc-based model and a node-oriented model, each designed to capture distinct structural aspects of the problem. We further introduce a preprocessing reduction method that eliminates redundant arcs, significantly enhancing computational performance and scalability. Both formulations are implemented using classical and quantum optimization approaches. In particular, the classical models are solved with Gurobi, whereas the quantum implementation is carried out on D-Wave’s LeapCQMHybrid platform, a hybrid quantum–classical environment that integrates quantum annealing with classical optimization techniques for constrained problem solving. Numerical experiments are conducted to validate the proposed formulations and the preprocessing reduction method. The analyses performed assess the structural properties of the two models, their computational behavior, and the impact of preprocessing on problem size and solution efficiency.
keywords:
Steiner Traveling Salesman Problem, Pickup and Delivery, Time Windows, Graph Reduction, QA, D-Wave
journal: a journal
[…]
[aff1]organization=Department of Mechanical, Energy and Management Engineering, University of Calabria, addressline=Via Pietro Bucci, city=Rende, postcode=87036, state=CS, country=Italy
[…]
[aff2]organization=TECNALIA, Basque Research and Technology Alliance (BRTA), addressline=Astondo Bidea, Edificio 700, city=Derio, postcode=48160, state=Bizkaia, country=Spain
1 Introduction
Routing problems are among the most central and extensively studied challenges in combinatorial optimization, due to their wide range of real-world applications in domains such as logistics, transportation and urban distribution. In recent years, the relevance of these problems has grown significantly, driven by the digital transformation of supply chains, the exponential rise of e-commerce, and the increasing demand for reliable, customized solutions. Effective route optimization not only reduces operational costs but also contributes to lower emissions, improved fleet utilization and higher service quality.
The Vehicle Routing Problem (VRP) is a fundamental combinatorial optimization problem in logistics and transportation. It focuses on determining the most efficient routes for a fleet of vehicles to deliver goods to a set of customers, starting and ending at one or more depots [ghiani2003real, braekers2016vehicle]. Among the most studied extensions of the VRP is the variant involving time windows, where each customer must be served within a specified time interval. This reflects realistic delivery scenarios, where users often select preferred delivery slots when placing an order. These temporal constraints add realism to the model but also significantly increase its computational complexity, as feasible solutions must simultaneously satisfy both spatial and temporal requirements [pugliese2017last, HAM20181, di2020trucks].
Another practically relevant variant is the VRP with pickup and deliveries (VRP-PD), in which each request involves a pair of locations, one for pickup and one for delivery, subject to precedence and capacity constraints. This model reflects the growing complexity of modern logistics systems, where vehicles must not only deliver but also collect items, such as in reverse logistics, product returns or shared mobility services. VRP-PD based models have a wide range of applications, including ride-sharing, e-commerce, hospital logistics and urban freight distribution [li2001metaheuristic, desaulniers2002vrp, ropke2006adaptive].
The simultaneous integration of time windows and pickup and delivery constraints results in a highly challenging class of routing problems, due to the combined presence of spatial, temporal and logistical dimensions. These problems are known to be NP-hard and their complexity grows rapidly with problem size, making them particularly difficult for both exact and heuristic classical methods.
Furthermore, the Steiner Traveling Salesman Problem (STSP) is a combinatorial optimization problem that generalizes both the Traveling Salesman Problem (TSP) and the Steiner Tree Problem (STP). It aims to find a minimum-cost route that visits a subset of required nodes, referred to as required nodes, in a graph at least once. The route may include optional Steiner nodes, non-required vertices that help reduce the overall cost and allows edges to be traversed multiple times, offering flexibility not present in the classical TSP.
From the TSP, the STSP inherits the objective of constructing an efficient tour. However, unlike the TSP, which requires visiting all nodes exactly once, the STSP only requires covering the required nodes and permits the reuse of nodes and edges. From the STP, it adopts the idea of connecting key nodes while using intermediate ones to minimize cost, though the STSP seeks a tour rather than a tree structure.
This hybrid structure makes the STSP particularly suitable for applications in network design, logistics, and transportation, where only a subset of critical points must be connected efficiently within an existing infrastructure. As with the TSP and STP, the STSP is NP-hard, and solving large instances exactly remains computationally challenging.
Despite its practical potential, extending the STSP to account for real-world constraints such as time windows and pickup and delivery has received limited attention in the literature, possibly due to the high computational complexity these extensions introduce.
In this work, we address a significant gap in the current literature by introducing the Steiner Traveling Salesman Problem with Time Windows and Pickup–and–Delivery (STSP-TWPD), a novel and practically motivated routing problem that, to the best of our knowledge, has not been explored previously. This new variant extends the classical STSP by incorporating both temporal feasibility requirements and pickup–delivery relations, thereby capturing key features of modern urban logistics scenarios.
To model this problem, we propose two mathematical formulations. The first, an Arc-Based Formulation (ABF), builds upon traditional STSP models by assigning routing variables to arcs and incorporating constraints to enforce service times, time-window consistency, and pickup–delivery precedence. The second, a Node-Based Formulation (NBF), introduces a different modeling perspective in which routing decisions are expressed at the node level.
The literature currently lacks formal definitions or dedicated analyses of the Steiner TSP with either time windows (STSP-TW) or pickup–and–delivery constraints (STSP-PD). We therefore provide the first explicit definitions of these variants and evaluate them within a unified modeling and experimental framework. This allows us to disentangle and quantify the individual effects of temporal and precedence constraints, while establishing new benchmark problems and a methodological baseline for future research on Steiner-based routing.
To support the solution of these problems, we introduce a classic preprocessing procedure that removes redundant arcs and reduces overall model complexity. This reduction step is designed to handle the additional structural intricacies introduced by time-window and pickup–delivery constraints in the STSP-TWPD, thereby improving tractability and computational efficiency.
We develop a reproducible benchmark generation procedure for STSP-TWPD instances. This includes the specification of rules for generating time windows, service durations, and pickup-delivery requests, providing a foundation for reproducible and meaningful experimental analysis.
In addition to the classical solution obtained with Gurobi, the study further examines a quantum–hybrid approach for solving the proposed problem.
In recent years, quantum computing has emerged as a novel paradigm for tackling combinatorial optimization problems that are computationally demanding for classical algorithms. Although it has shown potential in this field, handling complex constraint structures remains a major challenge for current quantum optimization approaches. Two main paradigms currently define the landscape of quantum computation: gate-based quantum computing and Quantum Annealing (QA).
Gate-based systems perform universal quantum computation by applying sequences of quantum logic gates to qubits, enabling algorithms such as Shor’s factorization [shor1994algorithms] and Grover’s search [grover1996fast], which theoretically provide exponential or quadratic speed-ups over classical methods. In practice, however, these algorithms require large-scale, fault-tolerant architectures with thousands of error-corrected qubits, resources that remain far beyond current technological capabilities [preskill2018quantum]. As a result, gate-based methods are mostly limited to proof-of-concept demonstrations and demand specialized expertise in quantum circuit design, state manipulation, and error correction, restricting their use in applied optimization contexts.
QA, on the other hand, represents a more specialized and technologically mature approach tailored for optimization. It exploits quantum tunneling and adiabatic evolution to search for the minimum of an energy landscape encoded as an Ising or Quadratic Unconstrained Binary Optimization (QUBO) model [glover2018tutorial, glover2020quantum]. This paradigm is particularly appealing because many combinatorial optimization problems can, in principle, be formulated as QUBOs, and modern hardware platforms such as D-Wave’s quantum annealers already allow the execution of optimization instances. QA has therefore been explored in a wide range of combinatorial optimization problems, including scheduling [venturelli2016job, perez2024solving, ciacco2025Educational], routing [holliday2025advanced], facility location [ciacco2026facility, malviya2023logistics], and packing [de2022hybrid, garcia2022comparative].
Despite these efforts, practical limitations persist. The associated energy landscapes often exhibit high ruggedness and numerous local minima, which can hinder the convergence of QA and variational algorithms [abbas2024challenges]. Recent analyses have also emphasized the gap between demonstrating feasibility of quantum formulations and achieving any meaningful notion of quantum advantage. In particular, Smith-Miles et al. [smith2025travelling] argue that current quantum optimization pipelines remain highly sensitive to modeling choices, penalty parameter tuning, hardware noise, and embedding constraints, and therefore exhibit no clear scaling benefits over state-of-the-art classical methods. Their conclusions reinforce the need for a cautious and honest interpretation of existing QA results, especially in routing problems where instance sizes remain small.
Although we are aware of the current technical limitations of quantum optimization, in this work we nevertheless explore the use of a Constrained Quadratic Model (CQM)-based approach to solve the problem. To this end, we employ the CQM framework implemented through D-Wave’s LeapCQMHybrid solver. Unlike traditional QUBO formulations, the CQM framework preserves explicit linear and quadratic constraints alongside the quadratic objective function, enabling feasibility to be enforced directly rather than through penalty terms. The hybrid quantum–classical architecture delegates constraint handling to a classical optimization layer while leveraging the quantum annealer to explore the combinatorial search space. This makes the CQM approach a more stable and expressive modeling environment for complex constrained optimization problems, reducing the reliance on manual penalty tuning and preserving the structural integrity of the original formulation.
The main contributions of this work are as follows:
  • 1.
    We introduce a novel and previously unaddressed variant of the STSP, the STSP-TWPD, which extends the STSP by incorporating time windows and pickup and delivery constraints;
  • 2.
    We propose two novel formulations for the STSP-TWPD: an arc-based model, ABF, and a node-based model, NBF;
  • 3.
    We formally define and analyze two additional Steiner-based routing variants (STSP-TW and STSP-PD), which, to the best of our knowledge, have not been previously addressed in the literature;
  • 4.
    We present a reduction method that enhances the models by eliminating unnecessary arcs and minimizing model complexity;
  • 5.
    We introduce a novel procedure for generating STSP-TWPD instances. In particular, we define a method for the creation of time windows, service times and customer demands for STSP-TWPD.
  • 6.
    We implement and solve the proposed models using classical optimization tools as well as D-Wave’s hybrid quantum–classical solver, assessing their applicability to instances of practical relevance.
The remaining of this paper is structured as follows. In Section 2, we review the scientific literature related to STSP, VRP-TWPD and quantum approaches for combinatorial optimization problems. Section 3 provides a detailed description of the problem under study, including its main characteristics and motivation. Section 6 reports the computational experiments, including the instance generation process, solver setup, and numerical analysis performed with both classical and quantum approaches. Section 7 summarizes the concluding remarks and prospects for future studies.
2 Related works
This section presents a brief overview of the current state of the art, aiming to position the proposed research within the broader scientific context. To the best of our knowledge, no prior work has directly addressed the STSP-TWPD, neither through classical optimization methods nor via quantum computing approaches such as QA.
To bridge this gap, the literature is reviewed along two key dimensions. First, we examine classical methods developed for the STSP and the VRP-TWPD, which form the methodological foundation for our work. Second, we discuss recent routing studies that employ quantum or hybrid quantum–classical optimization approaches, highlighting current progress and remaining limitations. The analysis provides insights into the evolution of relevant problem formulations and solution strategies, offering the conceptual and technical foundation for the novel contribution presented in this paper.
2.1 Classical approaches
2.1.1 Works on the STSP
The STSP was introduced by Cornuéjols et al. [cornuejols1985traveling] and Fleischmann [fleischmann1985cutting]. Cornuéjols et al. [cornuejols1985traveling] presented the Graphical Traveling Salesman Problem, a generalization of the TSP defined on arbitrary graphs, where nodes and edges can be traversed multiple times. The goal is to find a tour that visits all nodes at least once, allowing for repeated traversals. This formulation is particularly suited for structured environments with constrained connectivity, such as city maps or factory layouts. Fleischmann [fleischmann1985cutting] proposed an exact solution method for the STSP based on a cutting plane approach. His model only considers the subset of required nodes and relaxes the requirement of Hamiltonian cycles, allowing for shorter tours. The formulation introduces valid inequalities tailored to the STSP to strengthen the linear relaxation, enabling exact solution of instances on realistic road networks. Letchford et al. [letchford2013compact] introduced three compact formulations for the STSP, leveraging techniques developed for the classical TSP. These include the single-commodity flow, multi-commodity flow, and time-staged formulations. Computational results indicate that the multi-commodity formulation is especially powerful, capable of solving instances with over 200 nodes using standard branch-and-bound algorithms. Notably, it achieves the same linear programming lower bound as Fleischmann’s exponential model, representing a significant computational advancement. Rodríguez-Pereira et al. [rodriguez2019steiner] proposed a recent compact ILP formulation based on the observation that any optimal solution to the STSP traverses each edge at most twice. Their model employs two sets of binary variables to represent the first and second traversal of each edge, significantly reducing model size while preserving accuracy. This formulation enables exact solution of instances with up to 500 nodes, outperforming prior approaches. Interian and Ribeiro [interian2017grasp] developed a metaheuristic based on the Greedy Randomized Adaptive Search Procedure, enhanced with path-relinking and periodic restarts. Their approach incorporates a reduced 2-opt neighborhood, problem-specific greedy criteria, and local search intensification. The use of path-relinking allows structured exploration of elite solution paths, improving overall solution quality. Experimental results confirm its effectiveness on large, sparse STSP instances. Álvarez-Miranda et al. (2019) [alvarez2019note] proposed transforming the STSP into a standard TSP by computing the metric closure over the required nodes, where each edge represents the shortest-path distance in the original graph.
2.1.2 Works on the VRP-TWPDP
Dumas et al. (1991) [dumas1991pickup] proposed a foundational exact method for the VRP-TWPDP, incorporating vehicle capacity, pickup and delivery precedence, and time window constraints. Their approach is based on a column generation scheme applied to a set-partitioning formulation, where each column represents a feasible route. The subproblem, formulated as a resource-constrained shortest path, is solved via dynamic programming and provides strong lower bounds. Chang et al. (2003) [chang2003real] extended the VRP-TWPDP into a real-time environment, introducing a formulation where customer requests arrive dynamically and must be handled immediately. The authors proposed a heuristic composed of three main components: route construction, route improvement using an anytime algorithm, and a tabu search procedure. Parameters for the tabu search are optimized via Taguchi orthogonal arrays. Their results show that the method is efficient in reducing operational costs under real-time constraints. Both approaches provide solid foundations for practical vehicle routing under temporal and resource constraints. Fabri and Recht (2006) [fabri2006dynamic] proposed a dynamic extension of the VRP-TWPDP in which each customer request includes two separate time windows, one for pickup and one for delivery. Departing from the stretch-factor approach in [caramia2002routing], their model allows explicit time window constraints and vehicle waiting. The solution combines exact single-vehicle routing via dynamic programming and a fast heuristic to allocate requests across vehicles.
As shown in Table 1, the classical STSPVRP-PD (i.e., the variant without time-window constraints or paired pickup-and-delivery requirements) has been addressed at very large scales using both exact algorithms and sophisticated heuristics.
The problem considered in this work, however, presents a substantially richer structure. The addition of time windows combined with pickup-and-delivery operations introduces temporal feasibility constraints and precedence relationships that significantly alter the underlying combinatorial landscape. A related line of research is the VRP-TWPD, which represents the closest classical counterpart. Early exact methods, such as those by Dumas et al. [dumas1991pickup], already demonstrated how the interaction between time-window feasibility and pickup–delivery pairing severely restricts tractable instance sizes, with exact algorithms typically limited to a few dozen requests. Subsequent heuristic and metaheuristic studies achieve larger scales, up to a few hundred requests, yet consistently highlight the intrinsic difficulty of jointly managing precedence, time windows, and routing decisions.
Table 1: Comparison of the main contributions in the literature on the STSP and its variants and on VRP-TWPD and its related extensions. The table includes the following columns: Reference lists the original authors; STSP indicates whether the work addresses the classical STSP; TW specifies whether time windows are considered; PD identifies whether pickup-and-delivery constraints are included; Max size reports the largest instance solved (in number of nodes); Approach summarizes the nature of the method proposed (theoretical, exact, or heuristic).
Reference STSP TW PD Max size Approach
Cornuéjols et al. [cornuejols1985traveling] theoretical
Fleischmann [fleischmann1985cutting] 292 nodes exact
Letchford et al. [letchford2013compact] 200 nodes exact
Rodríguez-Pereira et al. [rodriguez2019steiner] 500 nodes exact
Interian and Ribeiro [interian2017grasp] 3353 nodes heuristic
Álvarez-Miranda et al. [alvarez2019note] 300 nodes exact
Dumas et al. [dumas1991pickup] 15 nodes exact
Chang et al. [chang2003real] 100 nodes heuristic
Fabri and Recht [fabri2006dynamic] 1000 nodes heuristic
2.2 Quantum Approaches
A broad range of combinatorial optimization problems have been explored through QA. For a comprehensive and structured discussion of existing QA applications, we refer the reader to the surveys by Ciacco et al. [ciacco2025review], Osaba et al. [osaba2022systematic], Yulianti and Surendro [yulianti2022annealing] and Muhamediyeva et al. [muhamediyeva2026optimization]. Ciacco et al. [ciacco2025review] offer a detailed examination of quantum algorithm design across multiple domains, including healthcare, finance, production planning, and logistics, together with a systematic classification of methodological frameworks. Osaba et al. [osaba2022systematic] present a unified and end-to-end review of eighteen years of research (2004–2021) at the intersection between quantum computing and routing problems, analysing 53 contributions and identifying emerging trends and open challenges. In addition, Yulianti and Surendro [yulianti2022annealing] provide a systematic review of QA implementations, mapping 229 studies across thirteen application areas and four methodological families. Their work emphasizes how QA research has expanded in parallel with the evolution of commercial quantum annealers, highlighting both performance gains and persistent hardware-driven limitations. Muhamediyeva et al. [muhamediyeva2026optimization] examines the practical limitations of both quantum annealing and gate-based frameworks, and provides an accessible discussion of the quantum approximation optimization and quantum alternating operator ansatz paradigms.
Within this wider landscape, the only quantum-based study explicitly addressing the STSP is presented by Ciacco et al. [ciacco2025steiner]. In their work, the authors reformulate the STSP as a QUBO model derived from an integer linear programming formulation, introducing a dedicated preprocessing phase aimed at removing redundant arcs and reducing the dimensionality of the problem before its quantum optimisation. The resulting model is evaluated using both the D-Wave LeapBQM simulator and a Quantum Processing Unit (QPU), demonstrating that the preprocessing step substantially enhances computational performance and contributes to obtaining higher-quality solutions.
Against this broader backdrop, we focus here on reviewing recent and particularly significant QA-based contributions that are most relevant to the problem addressed in this study, placing specific emphasis on routing-related formulations. In particular, our analysis concentrates on the most recent developments appearing after 2022, thereby capturing the latest progress in the use of QA for routing problems. This targeted selection reflects the current state of the art while avoiding the redundancy of replicating the exhaustive analyses already provided by the aforementioned surveys.
Tambunan et al. [tambunan2022quantum] investigate QA approaches to the VRP with Weighted Road Segments, formulating the problem as a QUBO and solving it on a D-Wave QPU. Le et al. [le2023quantum] address the Orienteering Problem, another routing-related variant, through a QUBO formulation executed on D-Wave systems. Sinno et al. [sinno2023performance] provide an extensive empirical evaluation of commercial QA platforms on the Capacitated VRP (CVRP), conducting over 30 hours of experiments using D-Wave hybrid solvers. Their findings reveal that while QPU time remains extremely small, solution quality degrades with increasing constraint density rather than with problem size alone. In particular, the authors highlighting how modelling choices, and especially the number and structure of constraints in the QUBO or CQM formulation, critically influence performance. Further contributions include the work of Mori and Furukawa [10.3389/fphy.2023.1129594], who introduce the Adjuster Routing Problem, a disaster-response VRP in which insurance companies must deploy large numbers of adjusters across affected areas. Their QUBO formulation incorporates multiple operational constraints, and numerical experiments on D-Wave demonstrate that both solution quality and computational effort are highly sensitive to parameter selection. Recent contributions have further extended QA to realistic logistics settings. Osaba et al. [osaba2024solving] consider advanced delivery settings involving heterogeneous vehicle fleets, delivery priorities, and dual capacity constraints (weight and volume). Their experiments on real-world instances using D-Wave’s hybrid CQM solver confirm the practical relevance of quantum–classical methods for operational logistics. Mario et al. [mario2024hybrid] propose two hybrid classical-quantum strategies for real-time route optimisation. Their methods decompose the CVRP into clustering and routing phases, using classical Fuzzy C-Means for capacitated clustering and QA for solving either CVRP or TSP subproblems. Osaba et al. [osaba2025package] investigate the VRP-TWPD, integrating realistic constraints such as simultaneous pickup–delivery operations, time-window restrictions, and vehicle-specific mobility limitations. Their results, obtained through D-Wave’s Hybrid CQM Leap framework on seven classes of instances, reinforce the feasibility and robustness of hybrid QA approaches for complex logistics and routing environments.
Table 2 summarizes the key characteristics of the QA-based studies discussed above, including the type of problem addressed, the largest instance size tested, and the quantum approach employed.
Table 2: Overview of key QA studies in routing-related combinatorial problems. The table includes the following columns: Reference, which lists the authors of each study; Problem, which specifies the combinatorial optimization problem addressed; Max size, which reports the largest instance solved; Quantum Approach, which summarizes the type of quantum method employed.
Reference Problem Max size Quantum Approach
Ciacco et al. [ciacco2025steiner] STSP 9 nodes Hybrid BQM D-Wave Leap
5 nodes QPU D-Wave
Tambunan et al. [tambunan2022quantum] VRP with Weighted Road Segments 5 vehicles QPU D-Wave
Le et al. [le2023quantum] Orienteering Problem 7 cities QPU D-Wave
Sinno et al. [sinno2023performance] CVRP 80 customers Hybrid CQM D-Wave Leap
Mori and Furukawa [10.3389/fphy.2023.1129594] Adjuster Routing Problem 20 adjusters Hybrid CQM D-Wave Leap
Osaba et al. [osaba2024solving] VRP 29 customers Hybrid CQM D-Wave Leap
Mario et al. [mario2024hybrid] CVRP 37 customers QPU D-Wave
Osaba et al. [osaba2025package] VRP-TWPD 24 nodes Hybrid CQM D-Wave Leap
3 Problem definition
We introduce the STSP-TWPD, that integrates two important real-world constraints: time windows and pickup and delivery requirements. In the STSP-TWPD, a single vehicle starts and ends at a central depot and must serve all required nodes, i.e. customers that may request either pickups or deliveries. Each required node is associated with a time window, defined by a lower and upper bound, during which service must begin, reflecting real-world constraints such as store opening hours or scheduled delivery slots. The pickup and delivery component imposes precedence constraints, whereby certain nodes correspond to pickup requests and others to delivery requests. Importantly, this differs from the standard VRP-PD, where goods are transported between specific origin and destination locations. In our application context, a customer cannot request both a pickup and a delivery service at the same time. Pickup must precede a delivery, and all operations must be carried out within the time window associated with each customer.
The objective of the STSP-TWPD is to determine a minimum route starting and ending at the depot that visits all required nodes within their respective time windows and satisfies all pickup and delivery precedence constraints. The route may traverse optional nodes to build a feasible and efficient path.
The STSP-TWPD is defined on a directed not complete graph $G=(V,A)$, where $V$ is the set of all nodes, including the depot (0) and customer nodes and $V_{r}\subseteq V$ is the subset of customer nodes (required nodes) that must be visited.
Each customer node $i\in V_{r}$ is associated with a time window $a_{i},b_{i}]$, where $a_{i}$ denotes the earliest time at which service at node $i$ can start and $b_{i}$ denotes the latest acceptable start time. Each customer node is also characterized by a service time $s_{i}$, which represents the time required to complete the pickup or delivery operation at node $i$. Furthermore, each node is associated with a demand value $d_{i}$, which quantifies the amount of goods handled at that location. A positive value of $d_{i}$ indicates a pickup, while a negative value indicates a delivery. The travel time associated with arc $k$ is denoted by $l_{k}$. Equivalently, the travel time to traverse the arc from node $i$ to node $j$ is denoted by $l_{ij}$. We assume a unit-speed vehicle, so that the travel time is numerically identical to the arc length.
The vehicle has a total capacity $Q$ and the route must ensure that the vehicle’s load never exceeds capacity $Q$ at any point.
We denote by $\delta^{+}(i)$ the set of arcs leaving node $i$ and by $\delta^{-}(i)$ the set of arcs entering node $i$.
The objective of the problem is to determine a route that minimizes the total traversal time, defined as the sum of the travel times associated with the selected arcs, while ensuring that all customer nodes in $V_{r}$ are visited within their respective time windows $a_{i},b_{i}]$. The solution must also respect vehicle capacity limits throughout the route and properly handle pickup and delivery operations.
The parameters used in the formulation of the STSP are summarized in Table 3.
Notation Description
$V$ set of all nodes (including depot and customers)
$V_{r}$ set of customer nodes (required nodes)
AA Set of arcs in the graph
$\delta^{+}(i)$ set of arcs leaving node $i$
$\delta^{-}(i)$ set of arcs entering node $i$
$a_{i}$ earliest time to begin service at node $i$
$b_{i}$ latest time to begin service at node $i$
$s_{i}$ service time at node $i$
$d_{i}$ demand at node $i$
$l_{ij}$ travel time between node $i$ and node $j$
$l_{k}$ travel time of arc $k\in A$
$Q$ vehicle capacity
Table 3: Notation used in the problem formulation.
4 Mathematical Formulation
The first formulation, referred to as the ABF, is an arc-based model, where decision variables are defined on the arcs of the graph to capture both vehicle movement and the sequence of visits. The second formulation, called the NBF, adopts a node, based approach. Here, variables are associated with nodes.
4.1 Arc-Based Formulation - ABF
The ABF models the routing decisions using binary variables associated with arcs and time steps. This formulation closely follows classical STSP representations, while extending them to incorporate time windows and pickup-and-delivery constraints. By indexing arcs over time, ABF enables a precise representation of routing sequences and transitions between nodes.
4.1.1 Decision Variables.
The variables used to define the model are:
$y_{k}^{t}$
[…]
$x_{i}^{t}$
[…]
$\tau_{i}^{t}$:
depart time at node $i$ during time step $t$
$q_{i}^{t}$:
capacity of the vehicle after servicing node $i$ at time $t$
The time index $t$ is defined over the interval $1\leq t\leq|A|$, where $|A|$ is the total number of arcs in the graph. This upper bound ensures that the model can capture any feasible route, including the longest possible path that traverses all arcs. By associating a time step to each potential arc traversal, the model preserves temporal consistency and allows for accurate sequencing of operations throughout the routing process.
4.1.2 Objective Function.
$\min\quad$ $\sum_{t=1}^{|A|}\sum_{k\in A}l_{k}\>y_{k}^{t}$ (1)
The objective function (1) minimizes the total travel time, calculated as the sum of the travel times $l_{k}$ of all arcs traversed across all time steps.
4.1.3 Routing Constraints.
$\sum_{k\in\delta^{+}(0)}y_{k}^{1}=1$ (2)
Constraints (2) ensure that the vehicle departs from the depot (node 0) exactly once at the beginning of the planning horizon ($t=1$), thereby establishing the starting point of the route.
$y_{k}^{1}=0,\quad\forall k\in A\setminus\delta^{+}(0)$ (3)
Constraints (3) enforce that, at time $t=1$, only arcs emanating from the depot can be used. All other arcs are restricted at this stage, preventing the vehicle from appearing anywhere other than the depot at the beginning.
$\sum_{t=1}^{|A|}\sum_{k\in\delta^{+}(0)}y_{k}^{t}=\sum_{t=1}^{|A|}\sum_{k\in\delta^{-}(0)}y_{k}^{t}$ (4)
Constraints (4) ensure flow conservation at the depot across the entire planning horizon: the number of departures from the depot must equal the number of returns. This guarantees that all vehicle routes are closed and that no vehicle is left outside the depot at the end.
$\sum_{k\in\delta^{-}(i)}y_{k}^{t}=\sum_{k\in\delta^{+}(i)}y_{k}^{t+1},\quad\forall i\in V\setminus\{0\},\;1\leq t<|A|-1$ (5)
Constraints (5) enforce route continuity for every node $i$ except the depot, at consecutive time steps. It balances incoming and outgoing arcs so that if the vehicle arrives at $i$ at time $t$, it must also depart at time $t+1$, avoiding breaks or gaps in the route.
4.1.4 Single service for customer.
$x_{i}^{t}\leq M\>\sum_{k\in\delta_{i}^{-}}y_{k}^{t},\quad\forall i\in V,\;1\leq t<|A|$ (6)
Constraints (6) establish a link between the service variable $x_{i}^{t}$ and the routing variable $y_{k}^{t}$. It states that a service at node $i$ and time $t$ (i.e., $x_{i}^{t}=1$) is only allowed if the vehicle actually arrives at node $i$ at that time $t$.
$\sum_{t=1}^{|A|}x_{i}^{t}=1,\quad\forall i\in V_{r}$ (7)
Constraints (7) require each customer node $i\in V_{r}$ to be visited exactly once within the planning horizon, ensuring all demands are served for problem feasibility.
4.1.5 Time Constraints.
$\tau_{j}^{t+1}$ $\geq\tau_{i}^{t}+s_{i}\>x_{j}^{t+1}+l_{k}-M(1-y_{k}^{t+1}),$
$\quad\forall k\in A:k\in\delta_{i}^{+},k\in\delta_{j}^{-},i\in V,j\in V,\;0\leq t<|A|-1$ (8)
Constraints (8) enforce temporal consistency between consecutive node visits. If the vehicle travels from node $i$ to node $j$ at time $t+1$ via arc $k$, the arrival time at $j$ must respect both the service time $s_{i}$ at $i$ and the travel time $l_{k}$. The Big-M term ensures the constraint is inactive when the arc is not used.
$\tau_{i}^{t}\leq M\>\sum_{k\in\delta_{i}^{-}}y_{k}^{t},\quad\forall i\in V,\;1\leq t<|A|$ (9)
Constraint (9) links the time variable $\tau_{i}^{t}$ to the routing decision variables $y_{k}^{t}$. It ensures that a time value is assigned to node $i$ at time $t$ only if the vehicle arrives at node $i$ via an incoming arc at that time. The use of the Big-M constant prevents the assignment of a non-zero time when no vehicle reaches the node, effectively deactivating the constraint in such cases.
$\tau_{i}^{t}\geq(a_{i}+s_{i})\;x_{i}^{t},\quad\forall i\in V_{r},\;1\leq t\leq|A|$ (10)
$\tau_{i}^{t}\leq(b_{i}+s_{i})\;x_{i}^{t}+M(1-x_{i}^{t}),\quad\forall i\in V_{r},\;1\leq t\leq|A|$ (11)
Constraints (10) and (11) impose time window constraints. If service at customer $i$ is scheduled at time $t$, the corresponding time variable $\tau_{i}^{t}$ must lie within the valid time window $a_{i}+s_{i},b_{i}+s_{i}]$. The Big-M term relaxes the upper bound when the node is not serviced.
4.1.6 Capacity Constraints.
$q_{0}^{0}=Q$ (12)
Constraint (12) sets the initial load of the vehicle at the depot equal to its full capacity $Q$. We assume that the vehicle starts empty and has its full capacity available for pickups or deliveries.
$q_{i}^{0}\leq 0,\quad\forall i\in V\setminus\{0\}$ (13)
Constraints (13) ensure that no node other than the depot has any available load capacity at time zero.
$q_{i}^{t}\leq Q\>\sum_{k\in\delta_{i}^{-}}y_{k}^{t},\quad\forall i\in V,\;1\leq t<|A|$ (14)
Constraints (14) couple the load variable $q_{i}^{t}$ with the routing decisions. They ensure that a non-zero load can be assigned to node $i$ at time $t$ only if the vehicle actually arrives at that node via some incoming arc. The load is also bounded above by the vehicle’s capacity $Q$, and the Big-M logic embedded via $Q$ deactivates the constraint when the node is not visited at that time.
$q_{j}^{t+1}$ $\geq q_{i}^{t}-d_{j}\>x_{j}^{t+1}-Q\left(1-y_{k}^{t+1}\right),$
$\quad\forall\>k\in A:k\in\delta_{i}^{+},k\in\delta_{j}^{-},i\in V,j\in V;\>0\leq t<|A|-1$ (15)
Constraints (15) govern the dynamic update of the vehicle’s load as it moves through the network over time. Specifically, if a transfer from node $i$ to node $j$ takes place via an active arc $k\in A$ (i.e., $k\in\delta_{i}^{+}$ and $k\in\delta_{j}^{-}$ with $y_{k}^{t+1}=1$), then the load at node $j$ at time $t+1$ ($q_{j}^{t+1}$) must be at least the load at node $i$ at time $t$ ($q_{i}^{t}$), plus the quantity $d_{j}$ associated with a pickup or delivery at node $j$, provided that $x_{j}^{t+1}=1$. The term $Q(1-y_{k}^{t+1})$ acts as a big-M relaxation, deactivating the constraint when arc $k$ is not used.
$\sum_{t=1}^{|A|}q_{0}^{t}=Q-\sum_{i\in V_{r}}d_{i}$ (16)
Constraint (16) ensures that the load carried by the vehicle upon returning to the depot reflects the total quantity delivered throughout the planning horizon. Specifically, this condition enforces that the cumulative load on arcs entering the depot equals the initial capacity $Q$ minus the total customer demand $\sum_{i\in V_{r}}d_{i}$. In practice, this constraint guarantees that no load is "lost" during the route and that the residual load at the final return to the depot matches the expected leftover capacity, preserving load consistency and avoiding artificial capacity loss.
4.1.7 Variable Domains.
$y_{k}^{t}\in\{0,1\},\quad\forall k\in A,\;1\leq t\leq|A|$ (17)
$x_{i}^{t}\in\{0,1\},\quad\forall i\in V_{r},\;1\leq t\leq|A|$ (18)
$\tau_{i}^{t}\geq 0,\quad\forall i\in V,\;0\leq t\leq|A|$ (19)
$0\leq q_{i}^{t}\leq Q,\quad\forall i\in V,\;0\leq t\leq|A|$ (20)
These domain constraints (17)–(20) define the feasible values of the decision variables. The binary variables $y_{k}^{t}$ and $x_{i}^{t}$ encode route choices and service decisions, while $\tau_{i}^{t}$ and $q_{i}^{t}$ represent continuous time and load states, bounded to preserve problem realism.
4.2 Node-Based Formulation - NBF
The NBF is a novel approach in which routing decisions are defined over the nodes forming the arcs of the network.
4.2.1 Decision Variables.
The variables used to define the model are:
$y_{ij}^{t}$
[…]
$x_{i}^{t}$
[…]
$\tau_{i}^{t}$:
depart time at node $i$ during time step $t$
$q_{i}^{t}$:
capacity of the vehicle after servicing node $i$ at time $t$
The time index $t$ is retained in the NBF to capture the temporal evolution of routing and service decisions. Although constraints are structured at the node level, time-indexed variables are necessary to model sequencing, time-window feasibility, and load propagation across consecutive steps. As defined in ABF formulation, the time index $t$ is in the range $1\leq t\leq|A|$, which guarantees sufficient temporal solution to represent any feasible route, including the longest one that traverses all arcs in the graph.
4.2.2 Objective Function.
$\min\sum_{(i,j)\in A}\sum_{t=1}^{|A|}l_{ij}\;y_{ij}^{t}$ (21)
The objective function (21) minimizes the total travel time by summing the travel time $l_{ij}$ of all arcs $(i,j)$ that are traversed.
4.2.3 Routing Constraints
$\sum_{j\in V:(0,j)\in A}y_{0j}^{1}=\sum_{t=1}^{|A|}\sum_{j\in V:(0,j)\in A}y_{0j}^{t}=1$ (22)
Constraints (22) ensure that the vehicle departs from the depot exactly once, and that this departure occurs at the beginning of the planning horizon ($t=1$).
$\sum_{j\in V:(i,j)\in A}y_{ij}^{1}=0,\quad\forall i\in V\setminus\{0\}$ (23)
Constraints (23) prevent any node other than the depot from initiating a route. At time step $t=1$, only the depot is allowed to have outgoing arcs. All other nodes must have zero outbound flow, ensuring a valid and centralized route start.
$\sum_{t=1}^{|A|}\sum_{j\in V:(0,j)\in A}y_{0j}^{t}=\sum_{t=1}^{|A|}\sum_{j\in V:(i,0)\in A}y_{i0}^{t}$ (24)
Constraint (24) ensures flow balance at the depot over the entire planning horizon. It states that the total number of departures from the depot (i.e., arcs of the form $(0,j)$ used at any time $t$) must equal the total number of returns to the depot (i.e., arcs of the form $(i,0)$). This guarantees that vehicle that leaves the depot also returns, preventing route fragments that do not end at the depot and ensuring closed routes.
$\sum_{j\in V:(i,j)\in A}y_{ij}^{t+1}=\sum_{j\in V:(j,i)\in A}y_{ji}^{t},\quad\forall i\in V\setminus\{0\},\;1\leq t<|A|$ (25)
Constraints (25) enforce flow conservation at each non-depot node across time. Specifically, the vehicle can only leave a node if it previously arrived there, enforcing temporal consistency and eliminating disconnected paths.
4.2.4 Single service for customer.
$x_{j}^{t}\leq M\;\sum_{i\in V:(i,j)\in A}y_{ij}^{t},\quad\forall j\in V,\;1\leq t<|A|$ (26)
Constraints (26) ensure a logical link between the service variable $x_{j}^{t}$ and the routing decision variables $y_{ij}^{t}$. They ensure that service at node $j$ at time $t$ can only occur if the vehicle actually arrives at node $j$ through some incoming arc $(i,j)$ at that time. The use of the Big-M constant allows the constraint to be satisfied when no incoming arc is used (i.e., when the node is not visited), thereby avoiding infeasibility.
$\sum_{t=1}^{|A|}x_{i}^{t}=1,\quad\forall i\in V_{r}$ (27)
Constraints (27) guarantee that each required node $i\in V_{r}$ is visited and serviced exactly once during the planning horizon. This requirement is fundamental to the feasibility of the solution, as it ensures full coverage of all customer demands.
4.2.5 Time constraints.
The following timing constraints match those in the ABF formulation and guarantee correct scheduling and adherence to time windows for vehicle visits along the route.
$\tau_{j}^{t+1}\geq\tau_{i}^{t}+s_{j}\>x_{j}^{t+1}+l_{ij}-M(1-y_{ij}^{t+1}),\quad\forall(i,j)\in A,\;0\leq t\leq|A|-1$ (28)
Constraints (28) ensure time feasibility between two consecutive nodes. If arc $(i,j)$ is traversed, the arrival time at node $j$ must be consistent with the service duration at $j$, the travel time and the departure from $i$. The Big-M term disables the constraint when arc $(i,j)$ is not used.
$\tau_{i}^{t}\leq M\>\sum_{j\in V:(ji)\in A}y_{ji}^{t},\quad\forall i\in V,\;1\leq t<|A|$ (29)
Constraints (29) link the arrival time variable $\tau_{i}^{t}$ to the routing decisions $y_{ji}^{t}$, ensuring time consistency with vehicle movements. Specifically, they enforce that a valid arrival time at node $i$ at time $t$ can only be assigned if the vehicle actually arrives at $i$ via some incoming arc $(j,i)$ at that time. The Big-M constant deactivates the constraint when no such arc is used, preventing $\tau_{i}^{t}$ from taking a non-zero value unless node $i$ is truly visited.
$\tau_{i}^{t}$ $\geq(a_{i}+s_{i})\;x_{i}^{t},\quad\forall i\in V_{r},\;1\leq t\leq|A|$ (30)
$\tau_{i}^{t}$ $\leq(b_{i}+s_{i})\;x_{i}^{t}+M(1-x_{i}^{t}),\quad\forall i\in V_{r},\;1\leq t\leq|A|$ (31)
Constraints (30) and (31) enforce that, if a required node $i$ is serviced at time $t$, the associated service must begin within its predefined time window $a_{i},b_{i}]$. The lower bound ensures the service does not start too early, while the upper bound prevents late arrivals. The inclusion of the service time $s_{i}$ within both bounds aligns the constraint with the full duration of the stop. The Big-M relaxation deactivates the constraint when no service occurs at that time.
4.2.6 Capacity Constraints.
The following capacity constraints are similar to those in the ABF formulation, ensure proper load management throughout the route.
$q_{0}^{0}=Q$ (32)
Constraint (32) sets the initial load of the vehicle at the depot to its full capacity $Q$. This defines the starting inventory available for deliveries at the beginning of the planning horizon.
$q_{i}^{0}\leq 0,\quad\forall i\in V\setminus\{0\}$ (33)
Constraints (33) enforces that only the depot can have a load at time zero.
$q_{i}^{t}\leq Q\;\sum_{j\in V:(j,i)\in A}y_{ji}^{t},\quad\forall i\in V,\;1\leq t<|A|$ (34)
Constraints (34) enforce consistency between the load variables and routing decisions. They allow the load $q_{i}^{t}$ at node $i$ and time $t$ to be positive only if the vehicle arrives at that node through an incoming arc. When no vehicle enters node $i$ at time $t$, the constraint ensures that $q_{i}^{t}=0$. The upper bound $Q$ also ensures that the carried load never exceeds the vehicle’s capacity.
$q_{j}^{t+1}$ $\geq q_{i}^{t}-d_{j}\>x_{j}^{t+1}-Q\left(1-y_{ij}^{t+1}\right),\quad\forall\>i\in V,\forall\>j\in V;0\leq t<|A|-1$ (35)
Constraints (35) ensure consistency in the propagation of load across the network over time. If there is a transfer from node $i$ to node $j$ via arc $(i,j)$ at time $t+1$, then the inventory at node $j$ at time $t+1$ ($q_{j}^{t+1}$) must be at least the inventory at node $i$ at time $t$ ($q_{i}^{t}$), plus any quantity $d_{j}$ delivered or picked up at $j$ (when $x_{j}^{t+1}=1$). The term $Q(1-y_{ij}^{t+1})$ ensures the constraint is only enforced when the arc is active.
$\sum_{t=1}^{|A|}q_{0}^{t}=Q-\sum_{i\in V_{r}}d_{i}$ (36)
Constraint (36) ensure that the total load associated with returns to the depot is consistent with the vehicle’s initial inventory and the deliveries made. It guarantees that the load remaining on the vehicle upon return equals the unused portion of its capacity, i.e., the initial load $Q$ minus the total customer demand $\sum_{i\in V_{r}}d_{i}$.
4.2.7 Variable Domains.
$y_{ij}^{t}\in\{0,1\},\quad\forall(i,j)\in A,\;1\leq t\leq|A|$ (37)
$x_{i}^{t}\in\{0,1\},\quad\forall i\in V,\;1\leq t\leq|A|$ (38)
$\tau_{i}^{t}\geq 0,\quad\forall i\in V,\;0\leq t\leq|A|$ (39)
$0\leq q_{i}^{t}\leq Q,\quad\forall i\in V,\;0\leq t\leq|A|$ (40)
Constraints (37) and (38) defines the decision variables binary. Constraints (39) ensure that times are non-negative for all nodes and time steps. Constraints (40) bound the vehicle load between 0 and the maximum capacity $Q$ at all nodes and time steps.
4.3 Considerations on the ABF and NBF Formulations
This section clarifies the distinctive features of the ABF and NBF formulations, examines their structural dimensionality, and discusses the role of linearization parameters, such as the Big-$M$ and capacity constants, that ensure model consistency and numerical stability.
4.3.1 Equivalence Between ABF and NBF Formulations
We prove that the ABF and the NBF are mathematically equivalent representations of the STSP-TWPD. Despite their structural differences, ABF defines binary arc-time variables $y_{k}^{t}$, while the NBF employs node-time variables $y_{ij}^{t}$, both models encode the same feasible space and lead to the same optimal solutions. We establish this equivalence by constructing bijective mappings between their variable sets, constraints, and objective functions.
Variable Correspondence
we define a one-to-one mapping between the variables used in the two models:
  • 1.
    Arc variables: $y_{ij}^{t}\equiv y_{k}^{t},\quad\text{where }k=(i,j)$. In both models, $y_{ij}^{t}$ (or $y_{k}^{t}$) denotes whether arc $(i,j)$ is traversed at time $t$. Thus, the arc-time variables are directly shared.
  • 2.
[…]
  • 3.
[…]
Objective Function Correspondence
the objective functions in both ABF (1) and NBF (21) are equivalent. Therefore, minimizing the total travel time in (1) is equivalent to minimizing the total travel time in (21).
Constraints Set Correspondence:
we verify that, for each category of constraints, the two models impose structurally identical restrictions. In particular:
  • 1.
    Routing and flow conservation: both models enforce that each route starts and ends at the depot and that flow is preserved throughout the route (ABF: Eq. (2)–(5), NBF: Eq. (22)–(25));
  • 2.
    Customer service: in both models, each request is served exactly once, and service variables are linked to routing decisions (ABF: Eq. (6) and (7); NBF: Eq. (26) and (27));
  • 3.
    Time and scheduling constraints: the models ensure time consistency across nodes and arcs and impose service within the given time windows, only when service is actually performed (ABF: Eq. (8)–(11); NBF: Eq. (28)–(31));
  • 4.
    Load and capacity: both formulations ensure that load propagation is consistent, delivery and pickup amounts are feasible and capacity limits are never exceeded (ABF: Eq. (12)–(16); NBF: Eq. (32)–(36));
  • 5.
    Variable domains: the domain definitions are the same in both models, including binary routing and service variables and bounded continuous time and load variables.
Model Dimensionality Comparison
Both the ABF and NBF have the same dimensionality when expressed in terms of the actual network size. Let $|V|$ be the number of nodes (including the depot), $|A|$ the number of directed arcs of the graph, $r=|V_{r}|$ the number of required nodes, and let $T$ denote the number of time layers used in the model (in our setting, $T=|A|$ as a safe upper bound). In the ABF, binary routing variables are defined for each arc–time pair, yielding $\Theta(|A|\cdot T)$ variables of type $y^{t}_{\cdot}$, plus $\Theta(r\cdot T)$ service variables $x_{i}^{t}$. Continuous variables for time and load, $\tau_{i}^{t}$ and $q_{i}^{t}$, contribute $\Theta(|V|\cdot T)$. The number of linear constraints is of the same orders, i.e., combinations of $\Theta(|A|\cdot T)$ and $\Theta(|V|\cdot T)$ terms due to routing/flow, timing, and capacity relations. The NBF employs the same indexing scheme in practice (arc–time for routing activation, node–time for service, time, and load), resulting in identical counts: $\Theta(|A|\cdot T)$ binary routing variables, $\Theta(r\cdot T)$ service binaries, and $\Theta(|V|\cdot T)$ continuous variables, with a comparable number of constraints. Hence, the two formulations are equivalent not only in feasibility and optimality but also in dimensionality.
The numerical results comparing the two formulations are presented in subsection 6.4.1, where we assess their performance and verify the correspondence of optimal solutions.
4.3.2 Linearization Parameters and Numerical Considerations
The constant $M$ is employed in both formulations to model time and capacity relations when the corresponding routing or service variables are not active. A similar reasoning applies to the vehicle capacity constant $Q$, which acts as a natural upper bound and plays a role analogous to that of $M$ in relaxing capacity-related constraints. While $Q$ has a clear physical interpretation, both constants function as upper-bounding parameters that indirectly affect the tightness of the relaxation and the numerical scaling of the model. However, the inclusion of such parameters inevitably increases the overall model complexity. Each constraint containing a Big-$M$ or $Q$ term introduces additional conditional logic that must be explicitly represented within the linear structure of the formulation. As a result, the number of constraints grows, and their coefficients span a wider numerical range, which can enlarge the branch-and-bound search tree and increase the solver’s computational burden. This effect becomes particularly evident in large-scale instances, where the accumulation of these upper-bounding constraints amplifies both the size of the constraint matrix and the number of potential branching points. To clarify their practical role within ABF and NBF models, the specific constraints in which these parameters are used are identified below. In the ABF formulation, $M$ appears in constraints (6), (9), and (11), while the capacity constant $Q$ is used in this role in constraints (14) and (15). Similarly, in the NBF formulation, $M$ appears in constraints (26), (29), and (31), whereas $Q$ is used in constraints (34) and (35). In all these cases, the Big-$M$ and $Q$ parameters allow the corresponding constraint to be automatically relaxed when the related routing or service decision variable is inactive, preserving linearity without introducing nonlinear or logical formulations.Similarly, in the NBF formulation, $M$ appears in constraints (26), (29), and (31), whereas $Q$ is used in this role in constraints (34) and (35). In all these cases, the Big-$M$ and $Q$ parameters allow the corresponding constraint to be automatically relaxed when the related routing or service decision variable is inactive. This mechanism is a standard linearization technique in mixed-integer linear programming (MILP), widely adopted to preserve model linearity and to avoid the introduction of logical or conditional expressions that would otherwise result in nonlinear formulations. Indeed, in the absence of this type of relaxation, several constraints would need to be expressed in a nonlinear form, involving products of binary and continuous variables.
For example, constraint (8) in the ABF model, which is written as:
$\tau_{j}^{t+1}\geq\tau_{i}^{t}+s_{i}x_{j}^{t+1}+l_{k}-M(1-y_{k}^{t+1}),$ (41)
would instead take the nonlinear form:
$\tau_{j}^{t+1}\geq(\tau_{i}^{t}+s_{i}x_{j}^{t+1}+l_{k})\;y_{k}^{t+1}.$ (42)
Similarly, for the load propagation constraint (15):
$q_{j}^{t+1}\geq q_{i}^{t}-d_{j}x_{j}^{t+1}-Q(1-y_{k}^{t+1})$ (43)
avoids the nonlinear form
$q_{j}^{t+1}\geq(q_{i}^{t}-d_{j}x_{j}^{t+1})y_{k}^{t+1},$ (44)
which would yield a non-convex feasible region.
Such nonlinearities would substantially increase computational complexity and reduce compatibility with classical MILP solvers as well as hybrid quantum optimization frameworks. Therefore, although the adoption of Big-$M$ and $Q$ parameters introduces known drawbacks in terms of numerical conditioning and model size, their inclusion is justified by the need to maintain a fully linear and computationally tractable formulation.
4.4 Problem Variants: STSP-PD and STSP-TW
The purpose of introducing these variants is twofold: to provide new benchmark instances for Steiner-based routing problems under temporal and operational constraints, and to isolate the computational effects of time-window and pickup-and-delivery conditions.
4.4.1 STSP-PD
To the best of our knowledge, the STSP-PD has not been previously formulated or analyzed as an independent variant of the STSP. This variant is derived directly from the ABF and NBF formulations presented in Subsection 4, with the sole modification of removing the time-window constraints (Eqs. (10)–(11) in the ABF and Eqs. (30)–(31) in the NBF). All other constraints remain unchanged from the full STSP-TWPD model. The STSP-PD variant thus provides a methodological baseline for evaluating how temporal feasibility constraints influence computational effort and model structure.
4.4.2 STSP-TW
Similarly, to the best of our knowledge, the STSP-TW has not been previously defined or investigated as an independent problem in the Steiner routing literature. We formally introduce this variant as a delivery-only extension of the STSP framework with time windows. While the STSP-TWPD jointly accounts for pickup-and-delivery precedence and temporal feasibility, the STSP-TW focuses exclusively on the temporal dimension, allowing us to isolate and examine the effects of time windows in the absence of pickup operations. From a modeling perspective, this requires adjustments to both instance generation and capacity initialization. In the standard STSP-TWPD setting, 50% of required nodes correspond to pickup requests (demands uniformly sampled from $\{10,\dots,20\}$) and 50% to deliveries (demands sampled from $\{-5,\dots,-1\}$), while all remaining nodes have zero demand. In the delivery-only case, all customer demands are instead drawn from $\{-5,\dots,-1\}$, effectively eliminating pickup operations. The vehicle, which in the STSP-TWPD starts empty with full capacity available ($q_{0}=Q$), must now begin fully loaded to perform only deliveries. Accordingly, the model initialization is modified as:
$q_{0}=0,$ (45)
and the flow conservation at the depot is updated to preserve consistency between total delivered and residual load:
$\sum_{t=1}^{|A|}q_{0}^{t}=Q+\sum_{i}d_{i}.$ (46)
These changes ensure that the STSP-TW remains feasible and structurally consistent with the underlying Steiner framework while capturing the pure delivery routing scenario.
5 Arcs Filtering and Graph Reduction Method
The Arcs Filtering and Graph Reduction (AFGR) method is a preprocessing technique applied to directed non-complete graphs, aimed at simplifying the original problem without losing essential information. Specifically, AFGR eliminates arcs that are either irrelevant or excessively costly (in terms of their travel time), while ensuring that the connectivity between required nodes, those that must be visited or serve a purpose in the solution, is fully preserved. By doing so, the graph is effectively reduced in size, which simplifies the underlying optimization problem. This simplification is fundamental for improving the efficiency of both the problem formulation and its subsequent solution process. By reducing the size of the arc set A, AFGR removes unnecessary complexity and focuses computational effort on the most promising parts of the search space. The resulting smaller and cleaner model leads to faster computations and more scalable algorithms. Although developed in the context of the STSP-TWPD, the AFGR procedure is general and can be readily adapted to other routing and network design problems that involve sparse or partially connected graphs, where preserving essential connectivity while reducing model complexity is crucial.
The method operates in several stages as described below:
  1. 1.
    Elimination of irrelevant arcs: we remove arcs that do not involve any required nodes. Specifically, only those arcs whose endpoints include at least one node from the required node set $V_{r}$ or the depot node 0 are retained.
  2. 2.
    Shortest-Path-Based Arc Filtering: for every pair of required nodes (i.e., all nodes in $V_{r}$ and the depot node 0), the algorithm computes shortest paths using Dijkstra’s algorithm. Any arc that participates in at least one such shortest path is considered essential and retained.
  3. 3.
    Retention of Useful Arcs: only the arcs involved in shortest paths between required node pairs are preserved. All other arcs are removed, since they are not part of any shortest route between critical nodes.
  4. 4.
    Eliminating Isolated Steiner Nodes: after arc filtering, the algorithm checks for isolated Steiner nodes, nodes not in $V_{r}$ and not incident to any remaining arc. These nodes are removed from the node set $V$.
  5. 5.
    Update graph structures: the remaining arcs are renumbered with sequential IDs to maintain a consistent indexing scheme. The graph’s internal data structures, including arc sets and the dictionaries for arc lengths, are updated accordingly.
To provide a clear and executable representation, the above steps are detailed in the pseudocode of 1.
Algorithm 1 AFGR
1:Input: $G=(V,A)$: directed graph, $V_{r}$: set of required nodes, 0: depot node
2:Output: $A^{\prime}$: filtered arc set
3:Step 1: Elimination of irrelevant arcs
[…]
5:Step 2: Shortest-Path-Based arc filtering
6:for each $u\in V_{r}\cup\{0\}$ do
7:  $(\text{dist},\text{prev})\leftarrow\text{Dijkstra}(G,u)$
8:  for each $v\in V_{r}\cup\{0\},v\neq u$ do
9:   if $v\in\text{prev}$ then
10:     Reconstruct path $u\rightarrow v$ using prev
11:     Mark all arcs in path as useful
12:   end if
13:  end for
14:end for
15:Step 3: Retention of useful arcs
[…]
17:Step 4: Eliminating isolated Steiner nodes
[…]
19:Remove any node $v\notin V_{r}\cup\{0\}$ that has no incident arc in $A^{\prime}$
20:Step 5: Update graph structures
21:Reindex arcs in $A^{\prime}$ sequentially
22:Update graph structures
23:return $V^{\prime},A^{\prime}$
A detailed quantitative analysis of the dimensionality reduction achieved by the AFGR procedure is presented in subsection (6.3). In that part of the paper, we evaluate how the application of AFGR affects the structural size of both mathematical formulations (ABF and NBF) by comparing the number of variables and constraints before and after preprocessing.
It is important to note that, although the AFGR procedure preserves full connectivity among required nodes, it does not provide a formal guarantee that the global optimal solution of the original problem is always retained after arc reduction. The method is therefore heuristic in nature, designed to improve computational tractability while maintaining the essential structural properties of the problem. Nevertheless, computational experiments indicate that the reduced graphs consistently yield optimal solutions across all tested instances.
6 Computational study
This section is divided into four main parts. The first part, Computational setup and solver configuration, describes the hardware and software environment, along with the solver details. The second part, Generation of instances, explains how the test instances for the experiments were created. The third part, Model dimensionality and computational complexity, analyzes the structural growth of the ABF and NBF formulations. Finally, the fourth part, Numerical results, presents and discusses the computational outcomes, highlighting the performance of the models and solvers.
6.1 Computational setup and solver configuration
All computational experiments are conducted on a Windows 10 machine equipped with an Intel processor (Family 6, Model 126, Stepping 5), featuring 4 physical cores (8 logical threads) and 15.6 GB of RAM. The implementation is developed in Python 3 using Jupyter Notebook. For exact optimization experiments, we employ Gurobi Optimizer version 11.0.1.
The CQM formulation is implemented using D-Wave’s Ocean SDK, specifically dimod version 0.12.18 and dwave-system version 1.28.0. All problem instances are solved through D-Wave’s hybrid solver service for CQMs, identified as hybrid_constrained_quadratic_model_version1p. We impose a time_limit of 5 seconds.
In practice, the CQM framework automatically converts each constraint into an energy-contributing term within the Hamiltonian by means of an internal penalty calibration mechanism. This automated process determines appropriate penalty scales that balance constraint satisfaction and objective minimization, ensuring that feasible solutions are prioritized without introducing numerical instability. As a result, users can focus on expressing the optimization model (defining variables, constraints, and objectives) without manually managing penalty coefficients. In our implementation, all decision variables are represented as binary or integer types, and both the objective and constraints are formulated using D-Wave’s dimod.ConstrainedQuadraticModel() interface, as documented in the official D-Wave Ocean SDK manual [DWaveCQMdocs]. The completed model is automatically compiled and submitted to the hybrid solver, which orchestrates the interplay between classical and quantum resources to enforce constraints and minimize the Hamiltonian energy.
From a methodological perspective, it is important to clarify the rationale for adopting the CQM framework instead of a standard QUBO formulation. A pure QUBO model represents an unconstrained binary quadratic problem in which all linear and mixed-integer constraints must be embedded into the objective function through quadratic penalty terms. This transformation introduces additional binary variables to encode continuous quantities (such as time and load) and requires large penalty coefficients to enforce feasibility.
The resulting energy landscape becomes highly ill-conditioned, and the appropriate penalty magnitudes must be empirically tuned to balance constraint satisfaction and objective optimization. Such manual calibration is problem-dependent and can strongly affect numerical stability, relaxation tightness, and convergence behaviour, issues that are further amplified in large-scale or hybrid quantum computing settings. In contrast, the CQM formulation allows constraints to be defined explicitly, maintaining the original linear structure of the MILP while avoiding artificial penalty weights and full binary encoding of continuous variables. This leads to a more stable and interpretable optimization model, in which feasibility is handled natively by the solver through an internal penalty-scaling mechanism.
We utilize the LeapCQMHybrid solver, a proprietary hybrid optimization tool developed by D-Wave Systemsa cloud-based framework designed to integrate classical and quantum computing resources within a unified optimization pipeline. As described in recent works on hybrid architectures [osaba2025d, bertuzzi2024evaluation, willsch2022benchmarking], the Hybrid Solver Service (HSS) follows a modular asynchronous workflow that combines the strengths of classical heuristics with quantum annealing-based exploration. From a structural perspective, each instance submitted to the hybrid solver is decomposed into multiple branches, which are executed in parallel on independent computational threads. Each branch contains two tightly coupled modules:
  • 1.
    Classical Module (CM): responsible for large-scale exploration of the solution space. The CM implements advanced metaheuristics, including local search, tabu search, and simulated annealing, to generate feasible candidate solutions and to construct reduced subproblems suitable for quantum processing;
  • 2.
    Quantum Module (QM): in charge of performing quantum-guided refinement. The QM executes quantum annealing cycles on the D-Wave Advantage2_system1.4 Quantum Processing Units (QPUs), which features 4596 qubits arranged in a Zephyr topology. Each quantum query corresponds to a partial embedding of the global CQM into the QPU’s architecture. The resulting quantum samples are analyzed and ranked according to their energy levels, then asynchronously sent back to the CM.
The CM and QM operate asynchronously, communicating through an adaptive feedback mechanism. The CM monitors convergence of its classical heuristics and invokes the QM whenever search stagnation is detected or diversification is required. This mechanism allows the quantum annealer to explore alternative regions of the energy landscape, providing probabilistic samples that can help the CM escape local minima. The asynchronous workflow also ensures that latency in one branch does not block the global progress of the solver as a major improvement over earlier synchronous hybrid frameworks. Figure 1 could schematically represent this interaction, showing how the CM initiates quantum subproblem queries, receives low-energy configurations from the QM, and updates its search trajectory based on quantum feedback.
Figure 1: General structure of D-Wave’s hybrid solvers, adapted from [osaba2025d]. Each branch includes a CM and a QM that interact asynchronously to refine candidate solutions.
The proportion of computation devoted to classical versus quantum phases is dynamically determined by the solver’s internal scheduler and is not user-controllable. According to D-Wave’s technical documentation [DWave2022performance] the frequency of quantum queries depends on problem size, constraint density, and classical progress metrics. In typical workloads, quantum queries account for a minority of total iterations but play a crucial exploratory role, guiding the search toward low-energy basins that are difficult to reach through purely classical heuristics. Although the internal configuration of the hybrid solver is proprietary, its design philosophy follows that of D-Wave’s hybrid architecture, where quantum and classical resources cooperate asynchronously. The classical component ensures scalability and constraint management, while the quantum component contributes through stochastic sampling based on quantum tunneling. This mechanism allows the solver to explore low-energy regions of the search space that are typically difficult to reach with classical heuristics alone. The use of the quantum module is therefore not aimed at achieving quantum speedup in a strict complexity-theoretic sense, but at improving diversification and refinement within the hybrid optimization loop. As evidenced in recent benchmark studies [osaba2025d, bertuzzi2024evaluation, willsch2022benchmarking], such hybridization can enhance convergence reliability and solution quality for large combinatorial problems that exceed the capacity of current QPUs.
It is therefore appropriate to regard the LeapCQMHybrid as a hybrid algorithm, where the quantum-guided component acts as an accelerator for diversification and refinement rather than as a full solver. The CM ensures scalability and feasibility management, while the QM provides stochastic sampling capabilities rooted in quantum tunneling effects. Together, these mechanisms form a synergistic loop that leverages both computational paradigms to achieve improved convergence speed and solution quality on combinatorial problems exceeding the current QPU capacity.
In our experiments, quantum queries are executed on D-Wave’s Advantage2_system1.4 QPU, the most advanced QPU architecture currently available. This system consists of 4596 qubits arranged according to a Zephyr connectivity graph. Solver parameters are kept at their default values, with the only tunable parameter, time_limit, left unchanged across all instances to ensure consistency and reproducibility. Due to the proprietary nature of LeapCQMHybrid, detailed information about its internal quantum routines, resource allocation, and the exact number of qubits used per instance is not disclosed. For further technical insights and practical applications, readers are referred to D-Wave’s official documentation and recent literature on hybrid quantum computing [osaba2025d, bertuzzi2024evaluation, willsch2022benchmarking].
6.2 Generation of instances
The node coordinates are generated by adapting the procedure proposed by Letchford et al. (2013)[letchford2013compact], a widely adopted benchmark method for constructing instances of the STSP. Specifically, we consider a set of $n$ vertices $V=\{v_{1},v_{2},\dots,v_{n}\}$ uniformly positioned along the circumference of a circle with fixed radius. Each vertex $v_{i}$ is assigned a Cartesian coordinate $(x_{i},y_{i})$ computed as:
$x_{i}=R\;\cos\left(\frac{2\pi i}{n}\right),\quad y_{i}=R\;\sin\left(\frac{2\pi i}{n}\right),$ (47)
where $R$ is the radius of the circle ($R=100$) and $i\in\{0,\dots,n-1\}$. From this layout, we consider all possible undirected edges between vertices (excluding the depot) and compute the Euclidean distance between endpoints as the base length of each arc. The candidate arcs are then sorted in ascending order of their length. Edges are incrementally added to the graph while ensuring the following constraints are satisfied:
  • 1.
    The new edge does not intersect any of the previously selected edges;
  • 2.
    The angle between any two adjacent edges incident to the same vertex is not less than $60^{\circ}$.
This selective insertion of edges yields a sparse, planar-like structure while preserving geometric properties inspired by real-world networks. To ensure connectivity, additional arcs are inserted where necessary. Specifically, for any vertex without incoming or outgoing arcs, at least one arc is artificially added to/from a randomly chosen adjacent vertex, subject to the same spatial consistency constraints. These additional arcs are assigned both a length equal to their Euclidean distance. Finally, all duplicate arcs, i.e., arcs defined on the same pair of nodes $(i,j)$, are removed to avoid redundancy and inconsistencies in the graph structure. In accordance with the instance generation procedure described by Letchford et al. [letchford2013compact], each node is independently designated as a required node with probability $p=0.7$, and one of the required nodes is randomly selected as the depot. This intermediate configuration, corresponding to approximately two-thirds of the nodes, yields instances of balanced density and computational complexity, in line with standard STSP benchmark practices.
For the parameters related to time windows, service times, demand, and vehicle capacity, we first carried out a literature review and identified Li and Lim (2001)[li2001metaheuristic] as a standard reference in the study of VRP-TWPDP.
We develop a custom instance generator specifically designed for the STSP-TWPD. The parameter generation follows a randomized procedure with a fixed seed to ensure reproducibility. The updated generation process is as follows:
  • 1.
    Time Windows ($a_{i}$, $b_{i}$): for each required node ($i\in V_{r}$), time windows are constructed based on both the shortest and longest path distances from the depot:
    • (a)
      The shortest paths $l_{\min}[i]$ are computed using Dijkstra’s algorithm and the longest paths $l_{\max}[i]$ via a Bellman-Ford’s strategy adapted for our graphs.
    • (b)
      Two weighted target values are used to introduce variability in the generation of time windows, ensuring a more diverse and realistic distribution:
      $l^{(1)}_{i}$ $=\alpha\cdot l_{\min}[i]+(1-\alpha)\cdot l_{\max}[i],$
      $l^{(2)}_{i}$ $=\beta\cdot l_{\min}[i]+(1-\beta)\cdot l_{\max}[i],$
      where $\alpha=0.7$ and $\beta=0.3$.
    • (c)
      The set of nodes $V_{r}$ is divided into two equal groups for generating early time windows:
      • i.
        For 50% of the nodes in $V_{r}$, the early window start time $a_{i}$ is generated as
        $a_{i}=l^{(1)}_{i}+\epsilon_{i},$
        where $\epsilon_{i}$ is a random value uniformly drawn from the interval $(0,\,0.5\cdot l^{(1)}_{i})$.
      • ii.
        For the remaining 50%, early windows are generated as
        $a_{i}=l^{(2)}_{i}+\epsilon_{i},$
        with $\epsilon_{i}$ is a random value taken from the interval $(0,\,0.5\cdot l^{(2)}_{i})$.
    • (d)
      The latest service start time $b_{i}$ is determined by adding a random value in $(l^{(2)}_{i},\;l^{(1)}_{i})$ to $a_{i}$, ensuring that $b_{i}\geq a_{i}$:
      $b_{i}=a_{i}+w_{i}$
    For non-required nodes ($i\in V\setminus V_{r}$), we fix $a_{i}=0$ and $b_{i}=10^{5}$.
  • 2.
    Service time ($s_{i}$): each required node receives a duration $s_{i}$ random in $\{5,\;10\}$. For all $i\in V\setminus V_{r}$, $s_{i}=0$.
  • 3.
    Demands ($d_{i}$): the nodes in the set $Vr$ are divided into two equal parts:
    • (a)
      50% of the nodes are assigned a positive demand $d_{i}\in[10,20]$, representing pickups.
    • (b)
      The remaining 50% are assigned a negative demand $d_{i}\in[-5,-1]$, representing deliveries.
    All nodes in $V\setminus V_{r}$ have zero demand.
  • 4.
    Vehicle Capacity ($Q$): the capacity is set to match the total absolute demand, with a small random offset:
    $Q=\sum_{i\in V}|d_{i}|+\theta,$
    where $\theta$ is random value in $\left(\min_{i}|d_{i}|,\;\max_{i}d_{i}\right)$.
    This promotes tight feasibility and favors the existence of feasible solutions.
  • 5.
    Big-M Constant ($M$): the constant $M$ is defined to upper-bound any feasible arrival time:
    $M=\max_{i\in V}b_{i}+\max_{i\in V}s_{i}+\max_{(i,j)\in A}l_{ij}$
    This expression combines three worst-case components:
    • (a)
      $\max_{i\in V}b_{i}$: the latest possible time window upper bound across all nodes;
    • (b)
      $\max_{i\in V}s_{i}$: the maximum service time at any node;
    • (c)
      $\max_{(i,j)\in A}l_{ij}$: the longest travel time between any pair of connected nodes.
    By summing these values, $M$ provides a safe upper bound on the arrival time at any node in any feasible route.
Table 4 summarizes the generation rules and notation.
Notation Description
$V$ Set of nodes
$V_{r}$ Subset of $V$ consisting of 70% of the nodes (required nodes)
AA Set of arcs derived from a sparse directed graph
$l_{\min}[i]$ Shortest path from depot to $i$ (Dijkstra’s algorithm)
$l_{\max}[i]$ Longest path from depot to $i$ (Bellman-Ford’s algorithm)
$a_{i}$ $\forall i\in V_{r}$: $a_{i}=t^{(1)}_{i}+\epsilon_{i}$ or $t^{(2)}_{i}+\epsilon_{i}$
$\forall i\in V\setminus V_{r}$: $a_{i}=0$
$b_{i}$ $\forall i\in V_{r}$: $b_{i}=a_{i}+w_{i}$, with $w_{i}$ random in $(l^{(2)}_{i},l^{(1)}_{i})$
$\forall i\in V\setminus V_{r}$: $b_{i}=100000$
$s_{i}$ $\forall i\in V_{r}$: $s_{i}$ random in $\{5,\dots,10\}$
$\forall i\in V\setminus V_{r}$: $s_{i}=0$
$d_{i}$ 50% of nodes in $V_{r}$: $d_{i}$ random in $\{10,\dots,20\}$ (pickups)
50% of nodes in $V_{r}$: $d_{i}$ random in $\{-5,\dots,-1\}$ (deliveries)
$\forall i\in V\setminus V_{r}$: $d_{i}=0$
$l_{ij}$ Euclidean distance between each pair of nodes $(i,j)$
$l_{k}$ Travel time of arc $k\in A$
$Q$ $Q=\sum_{i\in V}|d_{i}|+\theta$, $\theta$ random in $(\min_{i}|d_{i}|,\;\max_{i}d_{i})$
$M$ $M=\max_{i\in V}b_{i}+\max_{i\in V}s_{i}+\max_{(i,j)\in A}l_{ij}$
Table 4: Parameters and generation method for random STSP-TWPD instances.
6.3 Model Dimensionality and Computational Complexity
This section analyzes the dimensional growth and computational complexity of the proposed formulations, both in their original form and after applying the AFGR procedure. The analysis quantifies the evolution of the model size in terms of binary and continuous variables, as well as equality and inequality constraints, for increasing problem instances. The results highlight how AFGR effectively reduces model dimensionality, thus improving scalability and enabling the solution of larger instances. To assess the computational viability of the proposed approaches, we conduct an analysis of the two mathematical formulations introduced. The objective of this analysis is to compare the structural characteristics of the models in terms of the number of variables and constraints, which directly influence the scalability and solvability of the problem. The analysis is carried out both (ABF and NBF) with and without the application of the AFGR method, in order to highlight its impact on the structural size of the models. This study is performed across a range of instance sizes, allowing us to evaluate how the problem complexity evolves with increasing numbers of nodes and arcs.
Table 5: Detailed comparison of ABF complexity with and without AFGR method. For each instance size, we report: Bin: number of binary decision variables, Cont: number of continuous variables, Tot Vars: total number of variables (binary + continuous), $Eq:$ number of equality constraints, $\leq$: number of inequality constraints, $\geq$: number of inequality constraints, Tot C: total number of constraints. The columns labeled “With AFGR” show the reduced model size after applying the AFGR. The final columns indicate the percentage reduction in total variables and constraints.
$V$ Without AFGR With AFGR
% Red.
Vars
% Red.
Constr.
Bin Cont Tot Vars EqEq $\leq$ $\geq$ Tot C Bin Cont Tot Vars EqEq $\leq$ $\geq$ Tot C
4 77 64 141 31 108 84 223 77 64 141 31 108 84 223 0% 0%
5 150 110 260 51 184 160 395 150 110 260 51 184 160 395 0% 0%
6 280 180 460 84 313 308 705 280 180 460 84 313 308 705 0% 0%
7 494 280 774 132 500 551 1183 494 280 774 132 500 551 1183 0% 0%
8 825 416 1241 197 732 900 1829 494 280 774 132 500 551 1183 38% 35%
9 1312 594 1906 284 1064 1440 2788 494 280 774 133 519 570 1222 59% 56%
10 2000 820 2820 395 1489 2200 4084 825 416 1241 199 782 950 1931 56% 53%
11 2940 1100 4040 532 1970 3185 5687 825 416 1241 199 782 950 1931 69% 66%
12 4189 1440 5629 700 2607 4543 7850 1312 594 1906 286 1128 1504 2918 66% 63%
13 5810 1846 7656 901 3372 6300 10573 2000 820 2820 397 1569 2280 4246 63% 60%
14 7872 2324 10196 1138 4277 8528 13943 2940 1100 4040 535 2117 3332 5984 60% 57%
15 10450 2880 13330 1413 5239 11210 17862 2940 1100 4040 535 2117 3332 5984 70% 66%
16 13625 3520 17145 1731 6446 14606 22783 4189 1440 5629 703 2784 4720 8207 67% 64%
17 17484 4250 21734 2094 7828 18724 28646 5810 1846 7656 904 3582 6510 10996 65% 62%
18 22120 5076 27196 2504 9257 23520 35281 7872 2324 10196 1140 4441 8692 14273 63% 60%
19 27632 6004 33636 2966 11008 29359 43333 10450 2880 13330 1416 5524 11495 18435 60% 57%
20 34125 7040 41165 3482 12969 36225 52676 13625 3520 17145 1734 6773 14933 23440 58% 56%
Average Reduction: 47% 44%
Table 6: Detailed comparison of NBF complexity with and without AFGR method. For each instance size, we report: Bin: number of binary decision variables, Cont: number of continuous variables, Tot Vars: total number of variables (binary + continuous), $Eq:$ number of equality constraints, $\leq$: number of inequality constraints, $\geq$: number of inequality constraints, Tot C: total number of constraints. The columns labeled “With AFGR” show the reduced model size after applying the AFGR. The final columns indicate the percentage reduction in total variables and constraints.
$V$ Without AFGR With AFGR
% Red.
Vars
% Red.
Constr.
Bin Cont Tot Vars EqEq $\leq$ $\geq$ Tot C Bin Cont Tot Vars EqEq $\leq$ $\geq$ Tot C
4 77 64 141 29 108 84 221 77 64 141 29 108 84 221 0% 0%
5 150 110 260 48 184 160 392 150 110 260 48 184 160 392 0% 0%
6 280 180 460 79 313 308 700 280 180 460 79 313 308 700 0% 0%
7 494 280 774 124 500 551 1175 494 280 774 124 500 551 1175 0% 0%
8 825 416 1241 185 732 900 1817 494 280 774 124 500 551 1175 38% 35%
9 1312 594 1906 267 1064 1440 2771 494 280 774 125 519 570 1214 59% 56%
10 2000 820 2820 372 1489 2200 4061 825 416 1241 187 782 950 1919 56% 53%
11 2940 1100 4040 502 1970 3185 5657 825 416 1241 187 782 950 1919 69% 66%
12 4189 1440 5629 662 2607 4543 7812 1312 594 1906 269 1128 1504 2901 66% 63%
13 5810 1846 7656 854 3372 6300 10526 2000 820 2820 374 1569 2280 4223 63% 60%
14 7872 2324 10196 1081 4277 8528 13886 2940 1100 4040 505 2117 3332 5954 60% 57%
15 10450 2880 13330 1345 5239 11210 17794 2940 1100 4040 505 2117 3332 5954 70% 66%
16 13625 3520 17145 1651 6446 14606 22703 4189 1440 5629 665 2784 4720 8169 67% 64%
17 17484 4250 21734 2001 7828 18724 28553 5810 1846 7656 857 3582 6510 10949 65% 62%
18 22120 5076 27196 2397 9257 23520 35174 7872 2324 10196 1083 4441 8692 14216 63% 60%
19 27632 6004 33636 2844 11008 29359 43211 10450 2880 13330 1348 5524 11495 18367 60% 57%
20 34125 7040 41165 3344 12969 36225 52538 13625 3520 17145 1654 6773 14933 23360 58% 56%
Average Reduction: 47% 44%
Table 5 and Table 6 provide a detailed comparison of model complexity for increasing problem sizes ($V$), with and without the application of the AFGR method. The complexity is measured in terms of the total number of binary and continuous variables, as well as equality and inequality constraints.
Without the AFGR method, the size of both models increases rapidly as $|V|$ grows. Both the number of variables and constraints follow a super-linear, nearly exponential trend. For example, when $V$ increases from 4 to 20, the total number of variables rises from 141 to over 41000, and the total number of constraints grows from 223 to more than 52000. This growth illustrates the scalability challenges of the baseline model formulations, especially for large instances.
In contrast, applying the AFGR method significantly slows the growth in model size. While the number of variables and constraints still increases with $V$, the rate is considerably reduced. At $V=20$, for example, AFGR reduces the total number of variables from 41165 to 17145 and constraints from 52676 to 23440 in the ABF model. This corresponds to a 58% reduction in variables and a 56% reduction in constraints. Similar trends are observed in the NBF formulation, with consistent reductions beyond $V\geq 9$.
The benefits of AFGR are particularly evident for larger instance sizes:
  • 1.
    Variables: the original models exhibit near-exponential growth in the number of variables, whereas the AFGR-enhanced versions grow much more gradually, especially beyond $V=10$.
  • 2.
    Constraints: this is where AFGR delivers the most substantial impact, cutting constraint counts by more than half in many cases.
On average, AFGR achieves a 47% reduction in total variables and a 44% reduction in total constraints across all tested values of $V$. This confirms that AFGR not only reduces model size but also significantly improves scalability, enabling the solution of larger problem instances with greater computational efficiency.
(a) Total number of variables in the optimization model with and without AFGR as a function of $|V|$.
(b) Total number of constraints in the optimization model with and without AFGR as a function of $|V|$.
Figure 2: Comparison of the total number of variables (a) and constraints (b) in the optimization models with and without AFGR as a function of $|V|$. Orange lines represent the model with AFGR, while blue lines represent the model without AFGR. Since the ABF and NBF formulations exhibit identical trends in both variables and constraints (as shown in Table 5 and Table 6), the results are reported here in a single representative figure for clarity and conciseness.
Figure 2 illustrate the significant impact of incorporating AFGR on the model size. A consistent reduction is observed in both the total number of variables and constraints across all values of $V$ for both formulations. These results indicate that AFGR enables a more compact model formulation, which is potentially more computationally efficient.
6.4 Numerical results
The analysis of the results is organized into two main phases, each focusing on a distinct computational setting and on specific structural aspects of the problem.
  • 1.
    Performance analysis with a classical solver: we assess how the ABF and NBF formulations scale computationally across the three problem variants (STSP-TWPD, STSP-PD, and STSP-TW);
  • 2.
    Performance analysis with a quantum solver: we present and discuss the results obtained for the STSP-TWPD, STSP-PD, and STSP-TW using D-Wave’s hybrid quantum–classical solver. This phase evaluates the capability of the proposed ABF and NBF formulations to operate within a quantum optimization framework, assessing their performance in terms of solution quality, feasibility, and computational efficiency.
In each of the above settings, results are reported for both ABF and NBF models, with and without the AFGR procedure, to evaluate the performance impact of the proposed AFGR method. The outcomes include solution quality, constraint satisfaction and runtime analysis. Comparative tables and figures are provided to support the discussion and highlight the advantages and trade-offs among the different formulations and problem variants.
6.4.1 Performance analysis with a classical solver
We present a comprehensive computational analysis structured in four main stages. First, we examine the performance of the proposed formulations on the complete STSP-TWPD, assessing their scalability and the impact of the AFGR preprocessing method. Second and third, we extend the analysis to the two simplified variants, STSP-PD and STSP-TW, which are derived from the main formulation by selectively removing time-window or pickup-and-delivery constraints, respectively. Finally, we provide a comparative discussion among the three models (STSP-TWPD, STSP-PD, and STSP-TW), highlighting how their structural and computational characteristics evolve under different constraint configurations.
STSP-TWPD
We evaluate the computational performance of the STSP-TWPD model with two formulations (ABF and NBF). Table 7 reports the computational results obtained for the ABF and NBF formulations across instances of increasing size $V$. For each formulation, we present the objective value (OF), the computational time, and the optimality gap returned by Gurobi under a time limit of 5000 seconds.
Table 7: Comparison between the ABF and the NBF on instances of increasing size $V$. For each formulation, we report: OF, Time and GAP Opt. All experiments are performed with Gurobi using a time limit of 5000 seconds.
$V$ ABF NBF
OF Time (s) GAP Opt OF Time (s) GAP Opt
4 965.69 0.00 0% 965.69 0.02 0%
5 995.96 0.06 0% 995.96 0.03 0%
6 1019.62 0.08 0% 1019.62 0.09 0%
7 1401.16 3.13 0% 1401.16 3.52 0%
8 1429.73 6.84 0% 1429.73 5.89 0%
9 1299.21 12.05 0% 1299.21 5.97 0%
10 1620.41 742.86 0% 1620.41 177.25 0%
11 1559.06 2079.08 0% 1559.06 1019.55 0%
12 1324.50 136.00 0% 1324.50 68.70 0%
13 1062.23 49.47 0% 1062.23 31.92 0%
14 1350.57 3046.52 0% 1350.57 8207.86 0%
15 1333.68 2140.61 0% 1333.68 2279.67 0%
16 1498.74 15767.30 19% 1498.74 18817.78 16%
The results in Table 7 show that, for all instances where optimality is proven within the time limit, the ABF and NBF formulations return identical objective values. This empirical evidence supports the theoretical equivalence previously established in subsection 4.3.1. With respect to computational time, the two formulations display comparable behavior, with differences that are not systematic and can be attributed to solver-dependent branching decisions.
For instances with $V\geq 10$, both models exhibit a marked increase in computational effort, reflecting the growing complexity of the STSP–TWPD. The variability in running times across instances also suggests that performance is influenced not only by model size but by specific structural characteristics of each test case.
The instance with $V=16$ is particularly informative: although both formulations identify the same best integer solution, neither is able to prove optimality within the allotted time, resulting in nonzero optimality gaps. The difference between the reported gaps (19% for ABF and 16% for NBF) does not reflect discrepancies in solution quality but instead arises from the distinct tightness of their respective linear relaxation bounds. In other words, while the integer solutions coincide, the relaxations produce different lower bounds, which directly translate into the observed differences in optimality gaps.
We also assess the contribution of the AFGR method in improving the tractability and efficiency of the solution process. The experiments are conducted on instances with up to $V=10$ customers. Each model is tested in two configurations: with and without AFGR. All models are solved using Gurobi, with a time limit of 5 seconds for each instance. Table 8 report the results for the ABF and NBF formulations.
Table 8: ABF and NBF formulations of the STSP-TWPD model with and without the AFGR module. Each row corresponds to an instance with $V$ customers. For each formulation (ABF and NBF), results are shown both with and without AFGR. Metrics include: Vars: number of decision variables, FO: average objective function value across ten runs, Time [s]: average solution time in seconds, GAP opt: average optimality gap with respect to the best-known solution. The last two columns report the percentage variable reduction (% var red) for ABF and NBF.
$V$ ABF NBF
Without AFGR With AFGR % var red Without AFGR With AFGR % var red
Vars FO Time [s] GAP opt Vars FO Time [s] GAP opt Vars FO Time [s] GAP opt Vars FO Time [s] GAP opt
4 145 965,69 0,02 0% 145 965,69 0,00 0% 0% 145 965,69 0,02 0% 145 965,69 0,00 0% 0%
5 265 995,96 0,02 0% 265 995,96 0,02 0% 0% 265 995,96 0,00 0% 265 995,96 0,00 0% 0%
6 466 1019,62 0,06 0% 466 1019,62 0,08 0% 0% 466 1019,62 0,03 0% 466 1019,62 0,05 0% 0%
7 781 1401,16 2,95 0% 781 1401,16 3,78 0% 0% 781 1401,16 2,97 0% 781 1401,16 3,00 0% 0%
8 1249 1429,73 4,58 0% 781 1429,73 2,13 0% 37% 1249 1429,73 5,13 0% 841 1429,73 2,78 0% 33%
9 1915 1299,21 6,33 0% 781 1299,21 1,64 0% 59% 1915 1299,21 5,45 0% 901 1299,21 1,48 0% 53%
10 2830 1639,03 19,47 28% 1249 1620,41 16,84 0% 56% 2830 1678,47 19,41 43% 1405 1620,41 11,89 0% 50%
It is important to clarify that the reported optimality gap (GAP opt) values correspond to the Mixed-Integer Programming (MIP) gap returned by the classic solver Gurobi at termination. The MIP gap is a standard performance metric in mixed-integer optimization, defined as the relative difference between the best feasible solution (BestObj) and the best known lower bound (BestBound) at the time the solver stops:
$\text{MIP Gap (\%)}=100\times\frac{|\text{BestObj}-\text{BestBound}|}{|\text{BestObj}|}$ (48)
This metric quantifies how close the current solution is to proven optimality. A non-zero MIP gap typically indicates that the solver reached the predefined time limit before closing the gap entirely, rather than that the model provides a weak relaxation.
Both ABF and NBF formulations are able to find optimal solutions for instances up to $V=9$, with zero optimality gap. For $V=10$, both formulations show a marked increase in computational complexity. In the full ABF model (without AFGR), the solver returns a solution with a 28% optimality gap. The corresponding full NBF model performs worse, reaching a 43% gap. These results highlight the inherent difficulty of the problem at larger sizes and the need for strategies to improve tractability. Introducing AFGR significantly improves the performance of both formulations at larger instance sizes. In the ABF model, AFGR reduces the variable count by 56% at $V=10$, and enables the solver to close the optimality gap completely, achieving a 0% gap in the same time limit. A similar trend is observed in the NBF model, where AFGR achieves a 50% reduction in variables and reduces the optimality gap from 43% to 0%. In addition to improving solution quality, AFGR also reduces the average computational time in larger instances. For example, the average runtime for ABF at $V=10$ drops from 19.47s (without AFGR) to 16.84s (with AFGR), while for NBF it drops from 19.41s to 11.89s. This suggests that the simplified search space generated by AFGR improves solver efficiency without compromising solution quality. AFGR introduces no changes in small instances ($V\leq 7$), but leads to substantial reductions in the variables size as $V$ increases. The reduction reaches up to 56% for ABF and 50% for NBF at $V=10$, illustrating the effectiveness of AFGR in pruning infeasible or redundant decision variables in dense scenarios.
STSP-PD
We evaluate the performance of the STSP-PD using both the ABF and NBF formulations, tested with and without the AFGR module. Table 9 reports the detailed results for instances with up to $V=10$ customers.
Table 9: ABF and NBF formulations of the STSP-PD with and without the AFGR. Each row corresponds to an instance with $V$ customers. For each formulation (ABF and NBF), results are shown both with and without AFGR. Metrics include: Vars: number of decision variables, FO: average objective function value across ten runs, Time [s]: average solution time in seconds, GAP opt: average optimality gap with respect to the best-known solution. The last two columns report the percentage variable reduction (% var red) for ABF and NBF.
$V$ ABF NBF
Without AFGR With AFGR % var red Without AFGR With AFGR % var red
Vars FO Time [s] GAP opt Vars FO Time [s] GAP opt Vars FO Time [s] GAP opt Vars FO Time [s] GAP opt
4 145 965,69 0,00 0% 145 965,69 0,05 0% 0% 145 965,69 0,03 0% 145 965,69 0,05 0% 0%
5 265 995,96 0,00 0% 265 995,96 0,02 0% 0% 265 995,96 0,00 0% 265 995,96 0,02 0% 0%
6 466 1019,62 0,08 0% 466 1019,62 0,13 0% 0% 466 1019,62 0,13 0% 466 1019,62 0,09 0% 0%
7 781 1011,19 0,09 0% 781 1011,19 0,11 0% 0% 781 1011,19 0,08 0% 781 1011,19 0,05 0% 0%
8 1249 980,07 0,20 0% 781 980,07 0,13 0% 37% 1249 980,07 0,23 0% 841 980,07 0,20 0% 33%
9 1915 937,29 0,17 0% 781 937,29 0,09 0% 59% 1915 937,29 0,27 0% 901 937,29 0,20 0% 53%
10 2830 941,05 1,52 0% 1249 941,05 0,53 0% 56% 2830 941,05 1,48 0% 1405 941,05 0,81 0% 50%
All configurations solve to optimality within a negligible amount of time, with no impact from AFGR on either solution time or variable count. This confirms that, in the absence of time windows, the problem becomes significantly easier to solve, and both formulations remain equivalent in terms of performance and solution quality under these conditions.
As the number of customers increases, AFGR begins to significantly reduce the number of decision variables, up to 56% for ABF and 50% for NBF at $V=10$. Despite this reduction, all instances still reach optimality with a 0% gap, highlighting that removing time windows greatly simplifies the solution space. Moreover, AFGR contributes to lower solution times for larger instances. For instance, in the ABF formulation at $V=10$, the average time drops from 1.52s to 0.53s with AFGR; similarly, NBF drops from 1.48s to 0.81s. These results confirm that, even though the model complexity grows with $V$, the absence of temporal constraints makes the problem highly manageable from a computational standpoint.
When comparing these results with those obtained from the complete model including time windows (Table 8), the impact of the additional constraints becomes evident, particularly for instances with $V\geq 8$. In the complete model, solution times increase significantly, and the solver fails to close the optimality gap within the time limit for $V=10$ in both formulations without AFGR: the ABF model reaches a 28% gap, while the NBF model reaches 43%. In contrast, in the no-time-windows model, both formulations still achieve a 0% gap even at $V=10$, clearly indicating that time windows introduce considerable combinatorial complexity. AFGR proves to be a valuable component in both settings. However, its impact is more pronounced when time windows are included: in the complete model, applying AFGR restores optimality at $V=10$, while also halving the number of variables in both formulations. Nevertheless, solution times remain significantly higher compared to the no-time-windows scenario. For example, the ABF model with AFGR takes 16,84s at $V=10$, compared to 0,53s in the no-time-windows case.
Thus, time windows are an important factor in computational difficulty: their removal drastically reduces both solution times and model size. AFGR contributes meaningfully in both cases, but its utility is particularly evident when time windows are present, as it helps recover tractability in otherwise hard instances. The objective values remain identical across the two scenarios for each $V$, confirming that time windows do not alter the cost structure, but rather constrain the feasible set and increase solver effort.
STSP-TW
Table 10 presents the performance of the ABF and NBF formulations for the STSP-TW.
Table 10: ABF and NBF formulations of the STSP-TW with and without the AFGR. Each row corresponds to an instance with $V$ customers. For each formulation (ABF and NBF), results are shown both with and without AFGR. Metrics include: Vars: number of decision variables, FO: average objective function value across ten runs, Time [s]: average solution time in seconds, GAP opt: average optimality gap with respect to the best-known solution. The last two columns report the percentage variable reduction (% var red) for ABF and NBF.
$V$ ABF NBF
Without AFGR With AFGR % var red Without AFGR With AFGR % var red
Vars FO Time [s] GAP opt Vars FO Time [s] GAP opt Vars FO Time [s] GAP opt Vars FO Time [s] GAP opt
4 145 882,84 0,02 0% 145 882,84 0,02 0% 0% 145 882,84 0,02 0% 145 882,84 0,02 0% 0%
5 265 995,96 0,00 0% 265 995,96 0,02 0% 0% 265 995,96 0,05 0% 265 995,96 0,05 0% 0%
6 466 1019,62 0,02 0% 466 1019,62 0,02 0% 0% 466 1019,62 0,03 0% 466 1019,62 0,02 0% 0%
7 781 1028,38 0,27 0% 781 1028,38 0,17 0% 0% 781 1028,38 0,20 0% 781 1028,38 0,16 0% 0%
8 1249 1241,38 0,28 0% 781 1241,38 0,13 0% 37% 1249 1241,38 0,11 0% 781 1241,38 0,20 0% 37%
9 1915 1299,21 3,22 0% 781 1299,21 1,27 0% 59% 1915 1299,21 2,78 0% 781 1299,21 0,81 0% 59%
10 2830 1331,26 8,53 0% 1249 1331,26 2,28 0% 56% 2830 1331,26 5,36 0% 1249 1331,26 2,09 0% 56%
Both formulations yield identical performance in terms of solution time and optimality gap with and without AFGR. As the instance size increases ($V\geq 8$), AFGR starts to provide a significant reduction in model size, up to 59% fewer variables for both ABF and NBF at $V=9$. This reduction correlates with a consistent improvement in solution times, particularly for the ABF formulation. For instance, at $V=10$, the ABF with AFGR solves the instance in 2,28 seconds compared to 8,53 seconds without AFGR. In terms of objective function values, all configurations return the same results, indicating that AFGR does not compromise solution quality. Similarly, the optimality gap remains consistently at 0%, demonstrating the robustness of both formulations and the effectiveness of AFGR. Overall, the ABF and NBF formulations perform equivalently when AFGR is not used. However, when AFGR is applied, the ABF formulation tends to benefit more significantly in terms of computational time as problem size grows. A comparison between Tables 10 and 8 highlights the impact of including pickup operations in the STSP-TWPD model. The inclusion of pickup operations leads to a noticeable increase in the average objective function values, especially for larger instances. For example, at $V=10$, the objective function rises from 1331,26 (without pickup) to 1620,41 (with pickup) for ABF with AFGR, indicating a more complex scenario. Moreover, the solution times significantly increase when pickup operations are included. At $V=10$, the ABF formulation with AFGR requires 16,84 seconds with pickup, compared to just 2,28 seconds without. This trend is consistent across both formulations and highlights the additional computational burden introduced by modeling pickup constraints. Despite the increased complexity, the AFGR module remains effective. It consistently achieves substantial reductions in variable count, up to 59% for ABF and 56% for NBF in the no-pickup setting, and up to 56% and 50%, respectively, when pickup operations are included. This reduction translates into better scalability for larger instances.
Comparison of STSP-TWPD, STSP-PD and STSP-TW
The main objective of this set of experiments is to understand the role of the different constraints that define the STSP-TWPD problem. Since we propose STSP-TWPD as the most realistic variant, we compare it against its simplified versions (STSP-PD and STSP-TW) in order to assess how the solution changes when either time windows or pickup-and-delivery constraints are removed. We compare the performance of the three models, STSP-TWPD, STSP-PD and STSP-TW, by analyzing the routes generated for different instances. For this purpose, we consider instances with $V=7$, $V=8$, $V=9$, and $V=10$, which represent the largest and most informative instances available in our dataset. For each instance and each model, we visualize the optimal routes and analyze their characteristics. This visual inspection allows us to understand how time windows and pickup-delivery constraints influence the route structure. Moreover, we provide a quantitative comparison by plotting the objective function values for all models across the different instances. This plot highlights how the models perform relative to each other as the number of nodes increases. The figures below show the routes generated by each model for each instance size. Each subfigure corresponds to a specific model and instance size, and all are followed by a joint discussion comparing the structural differences among the models.
Figure 3: Comparison of optimal routes generated by the STSP-TWPD, STSP-PD and STSP-TW models for $V=7$.
Figure 3 illustrates the optimal routes generated by the three variants of the problem: STSP-TWPD, STSP-PD, and STSP-TW, for an instance with $V=7$ nodes. The visual differences in the route structures provide clear insight into the distinct effects of the modeled constraints. The route generated by the STSP-TWPD model (which incorporates both time windows and pickup-and-delivery constraints) appears highly structured and operationally realistic. The solution exhibits a tightly controlled path that respects both the precedence between pickup and delivery pairs and the feasibility of visiting nodes within their respective time windows. Arcs are shorter, the sequence is more intuitive, and the route avoids unnecessary detours, reflecting a careful optimization process under strict feasibility conditions. In contrast, the STSP-PD model omits time window constraints. While it continues to respect pickup-before-delivery precedence, the lack of temporal feasibility checks introduces greater spatial flexibility. This is reflected visually by more dispersed routing, with longer arcs and a less compact structure. Although the model adheres to precedence rules, the absence of time constraints can lead to suboptimal arrival times in practical settings. The STSP-TW model, on the other hand, enforces time windows but disregards pickup-and-delivery dependencies. As a result, the route structure prioritizes visiting nodes within their allowed time intervals, but without enforcing any service order between related pickup and delivery nodes. This leads to the most irregular and unstructured routing among the three models. Nodes may be visited in a sequence that is infeasible from a logistical standpoint, though still valid temporally.
Figure 4: Comparison of optimal routes generated by the STSP-TWPD, STSP-PD, and STSP-TW models for $V=8$.
Figure 4 presents the optimal routes derived from the STSP-TWPD, STSP-PD, and STSP-TW models for an instance with $V=8$ nodes. The differences in routing patterns highlight the impact of the specific constraints imposed by each model. The STSP-TWPD solution, which integrates both time windows and pickup-and-delivery constraints, displays a well-organized and realistic route. The path maintains strict adherence to pickup-before-delivery precedence and ensures all visits occur within designated time windows. The result is a compact, efficient route that minimizes unnecessary movement. In the STSP-PD model, time windows are excluded, leaving only the precedence relationships between pickup and delivery pairs. This relaxation introduces more spatial freedom, leading to visibly longer and more dispersed arcs. While precedence is preserved, the lack of temporal constraints may reduce practical feasibility in time-sensitive scenarios. The STSP-TW model enforces time windows but disregards the service order between pickup and delivery nodes. Consequently, while the route complies with timing requirements, it may violate logical delivery flows. This often results in suboptimal or counterintuitive sequences, underscoring the limitations of omitting precedence relationships.
Figure 5: Comparison of optimal routes generated by the STSP-TWPD, STSP-PD, and STSP-TW models for $V=9$.
Figure 5 shows the optimal routing solutions for the STSP-TWPD, STSP-PD, and STSP-TW models on an instance with $V=9$ nodes. As the problem size increases, the differences introduced by each model’s constraints become even more evident in the routing structures. The STSP-TWPD model, which simultaneously considers both time windows and pickup-and-delivery constraints, generates the most coherent and operationally sound route. The red arcs form a compact and efficient sequence, carefully obeying both temporal feasibility and precedence relations. The path avoids detours and demonstrates strong spatial and logical organization. The STSP-PD model relaxes the time window requirements while preserving pickup-before-delivery precedence. The resulting route reflects more flexibility in arc selection, leading to longer connections and a less compact structure. Although logically correct, the solution may be inefficient in time-sensitive operations due to its disregard for temporal constraints. The STSP-TW model enforces time windows but omits pickup-and-delivery relationships. This can be seen in the irregularity of the node visitation order, while all nodes are served within their respective time windows, there is no guarantee that pickups precede their associated deliveries. This leads to a route that may be temporally valid but infeasible from a logistics perspective.
Figure 6: Comparison of optimal routes generated by the STSP-TWPD, STSP-PD, and STSP-TW models for $V=10$.
Figure 6 presents the optimal routing solutions produced by the STSP-TWPD, STSP-PD, and STSP-TW models for an instance with $V=10$ nodes. As problem complexity increases, the distinctions introduced by each set of constraints become even more prominent. In the STSP-TWPD model, which integrates both time window and pickup-and-delivery constraints, the route is highly structured and operationally consistent. The solution respects the required precedence relationships and ensures that each node is visited within its feasible time window. The resulting route is compact, with efficient use of arcs and minimal detours, reflecting a balanced solution in terms of both spatial and temporal feasibility. In the STSP-PD model, which omits time windows, the route remains logically sound in terms of pickup-before-delivery ordering but is more spatially dispersed. The lack of time constraints allows for longer arcs and less compactness, which may lead to inefficiencies in real-world applications where timing is critical. The STSP-TW model, conversely, enforces time windows but ignores pickup-and-delivery precedence. The resulting route fulfills all temporal constraints but may violate service logic, visiting deliveries before their associated pickups. This leads to a less practical route, even if temporally valid, and highlights the limitations of ignoring precedence in time-constrained settings.
Figure 7: Objective function values for STSP-TWPD, STSP-PD, and STSP-TW models as a function of $V$.
Figure 7 clearly illustrates the evolution of the objective function values as the number of nodes $V$ increases. While all three models show similar performance for small problem sizes ($V\leq 6$), their behavior diverges significantly for larger instances. The STSP-TWPD model exhibits a sharply increasing trend, particularly for $V>6$. This reflects the compounded complexity introduced by simultaneously enforcing both time windows and pickup-and-delivery precedence. These combined constraints drastically reduce the solution space, often forcing the solver to take more costly routes in order to satisfy both feasibility requirements. In contrast, the STSP-PD model consistently yields much lower objective values beyond $V=6$. Since it does not impose any time window constraints, the model benefits from a more flexible temporal domain. This enables shorter and more direct routes, even while still respecting pickup-before-delivery precedence. The relative flatness of the red curve highlights the reduced sensitivity of STSP-PD to instance size compared to the more constrained models. The STSP-TW model occupies an intermediate position between the two extremes. It enforces time windows but ignores pickup-and-delivery ordering. As a result, the green curve often lies between the STSP-PD and STSP-TWPD lines. Although it adheres to temporal feasibility, the absence of service order constraints can lead to logically infeasible solutions, such as delivering before picking up, that reduce the objective cost but render the route impractical in real-world logistics settings. This comparison highlights the trade-off between solution quality (in terms of objective cost) and operational feasibility. While relaxing constraints such as time windows or precedence can yield cheaper solutions, only the STSP-TWPD model guarantees operationally realistic results under strict feasibility requirements.
6.4.2 Performance Analysis with a quantum solver
The objective is to assess the feasibility of expressing and solving the proposed formulations within a constrained quantum optimization framework. The experiments are intended as a proof of concept, providing insight into how the STSP-TWPD, and also STSP-PD and STSP-TW variants, can be represented and tackled through emerging hybrid quantum–classical architectures, specifically D-Wave’s LeapCQMHybrid platform. This part of the work is structured into three main phases. First, we analyze the results obtained for the STSP-TWPD under both the ABF and NBF formulations. Second, we examine a relaxed version of the problem, STSP-PD, in which the time window constraints are removed. Finally, we consider a further simplified scenario, STSP-TW, where only deliveries are allowed (i.e., no pickups), in order to isolate and better understand the individual impact of each modeling constraint on the overall computational complexity. All tables share the same structure. Each row corresponds to a specific problem size, indicated by $V$, which denotes the number of nodes in the instance. The columns are organized into two main groups, comparing the performance with and without the use of AFGR. For both configurations, the reported metrics include: the average and standard deviation of the objective function value (OF Avg and OF Std), the percentage of runs leading to a feasible or optimal solution over ten repetitions per instance (% Solved), as well as the average and standard deviation of computation time (Time Avg and Time Std).
We have conducted 10 tests for each instance in order to evaluate the performance and robustness of the model.
The upcoming paragraphs detail the numerical outcomes of each problem variant and discuss the impact of AFGR on solution quality and computational performance across varying instance sizes and data structures.
STSP-TWPD
This part of the work discusses the evaluation of solved problem instances. We compare the effectiveness of both the ABF and NBF formulations, examining cases with and without the implementation of the AFGR technique.
Table 11: Results for the STSP-TWPD using the ABF formulation, comparing performance with and without the AFGR method. Each row corresponds to a different instance size, defined by the number of vehicles $V$ and includes: OF opt: optimal OF value obtained with Gurobi, OF Avg: average OF value across ten runs, OF Std: standard deviation of the OF value, GAP: average optimality gap, calculated as the percentage deviation of OF Avg from OF opt, % Solved: percentage of instances solved successfully within the time limit, Time Avg: average solution time in seconds, Time Std: standard deviation of the solution time. A value of “–” indicates that no valid solution was obtained in any of the ten runs.
$V$ OFoptOF\>opt Without AFGR With AFGR
OF Avg OF Std GAP opt % Solved Time Avg Time Std OF Avg OF Std GAP opt % Solved Time Avg Time Std
4 965,69 965,69 0,00 0% 100% 11,78 0,28 965,69 0,00 0% 100% 11,99 1,10
5 995,96 1387,73 124,50 28% 100% 12,33 0,26 1333,44 183,33 25% 100% 12,46 0,59
6 1019,62 1917,22 130,94 47% 30% 13,55 0,60 1879,42 51,44 46% 60% 13,39 0,43
7 1401,16 - - - 0% 15,12 0,32 2715,55 0,00 48% 10% 15,73 1,77
8 1429,73 - - - 0% 17,48 0,93 2673,82 111,65 47% 20% 14,87 0,25
9 1299,21 - - - 0% 20,49 0,58 2349,56 0,00 45% 10% 15,32 0,76
10 4221,73 - - - 0% 24,90 0,60 - - - 0% 17,43 0,52
Table 12: Results for the STSP-TWPD using the NBF formulation, presented both with and without the AFGR method. Each row corresponds to a different instance size, defined by the number of vehicles $V$, and includes: OF opt: optimal OF value obtained with Gurobi, OF Avg: average OF value across ten runs, OF Std: standard deviation of the OF value, GAP: average optimality gap, calculated as the percentage deviation of OF Avg from OF opt, % Solved: percentage of instances solved successfully within the time limit, Time Avg: average solution time in seconds, Time Std: standard deviation of the solution time. A value of “–” indicates that no valid solution was obtained in any of the ten runs.
$V$ OFoptOF\>opt Without AFGR With AFGR
OF Avg OF Std GAP opt % Solved Time Avg Time Std OF Avg OF Std GAP opt % Solved Time Avg Time Std
4 965,6854 965,69 0,00 0% 100% 12,05 0,45 965,69 0,00 0% 100% 11,24 2,52
5 995,9593 1390,17 248,32 28% 90% 13,41 1,36 1371,49 218,9 27% 100% 16,43 11,89
6 1019,615 2079,42 183,23 51% 60% 15,00 2,62 1928,23 161,03 47% 60% 16,07 7,41
7 1401,163 - - - 0% 19,66 11,79 2273,17 251,44 38% 20% 16,77 3,91
8 1429,732 - - - 0% 21,37 8,46 2616,4 0,00 45% 10% 15,81 1,06
9 1299,211 - - - 0% 22,33 4,85 2508,5 119,37 48% 30% 15,95 1,38
10 4221,731 - - - 0% 26,48 6,23 - - - 0% 19,53 3,81
The experimental results reported in Tables 11 and 12 indicate that the AFGR technique contributes to improving the scalability and numerical stability of both the ABF and NBF formulations for the STSP-TWPD.
For the ABF formulation, AFGR consistently improves the solution quality and feasibility across all instance sizes, as shown in Table 11. On small instances ($V=4$ and $V=5$), both configurations solve 100% of the problems; however, AFGR provides a slightly better average GAP and more stable solution times. As the problem size increases ($V\geq 6$), the benefit of AFGR becomes even more pronounced: without AFGR, no feasible solutions are found for $V=7$ to $V=10$, while with AFGR, feasible solutions are obtained for $V=7$, $V=8$, and $V=9$, with solving rates up to 20%. Notably, the GAP with AFGR remains close to or slightly lower than in the baseline formulation when a solution is available (e.g., 46–48% for V=6V=6V=9V=9), whereas the baseline often fails entirely. Additionally, average solution times are comparable or even lower with AFGR for larger instances (e.g., $V=9$: 15,32s vs 20,49s), despite the increased complexity of the feasible repair mechanism. This demonstrates that AFGR enhances both efficiency and effectiveness.
Similarly, the NBF formulation benefits from AFGR across the full range of tested sizes. Table 12 reports the results for the STSP-TWPD solved using the NBF formulation, both with and without the AFGR method. While both variants perform similarly for small instances ($V=4$ to $V=6$), AFGR slightly reduces the optimality gap (e.g., 47% vs 51% for $V=6$), improves the percentage of solved instances (e.g., 100% vs 90% for $V=5$), and maintains acceptable solution times. For larger sizes ($V\geq 7$), the baseline fails to find feasible solutions, whereas AFGR achieves partial solving success (e.g., 20–30% for V=7V=7V=9V=9) with significantly improved GAP values (e.g., 38–48%).
Overall, these results confirm that the integration of AFGR contributes to improving both the feasibility and quality of solutions in both ABF and NBF formulations. AFGR increases the percentage of solved instances, generally reduces the optimality GAP and often leads to comparable or faster solution times.
STSP-PD
This part of the work analyzes STSP-PD problem. At this stage, we focus solely on the ABF formulation, rather than testing both ABF and NBF, as their equivalence has already been established in subsection (4.3.1). Scenarios with and without the AFGR mechanism are considered.
Table 13: Results of the STSP-PD problem, with and without AFGR. Each row corresponds to a different instance size, defined by the number of vehicles $V$, and includes: OF opt: optimal OF value obtained with Gurobi, OF Avg: average OF value across ten runs, OF Std: standard deviation of the OF value, GAP: average optimality gap, calculated as the percentage deviation of OF Avg from OF opt, Time Std: standard deviation of the solution time. A value of “–” indicates that no valid solution was obtained in any of the ten runs.
$V$ OFoptOF\>opt Without AFGR With AFGR
OF Avg OF Std GAP opt % Solved Time Avg Time Std OF Avg OF Std GAP opt % Solved Time Avg Time Std
4 965,69 965,69 0,00 0% 100% 13,67 5,77 965,69 0,00 0% 100% 12,31 1,18
5 995,96 1198,58 180,58 17% 100% 12,83 1,36 1100,7 141,43 10% 100% 12,65 1,51
6 1019,62 1715,50 290,64 41% 100% 14,21 1,65 1844,78 79,30 45% 100% 15,72 4,95
7 1011,19 2523,90 166,58 60% 100% 17,62 8,59 2517,81 248,58 60% 100% 16,26 4,70
8 980,07 3185,47 60,52 69% 30% 18,13 3,68 2614,67 165,32 63% 100% 15,10 1,31
9 937,29 - - - 0% 26,21 6,63 2419,53 292,51 61% 80% 16,95 5,78
10 941,05 - - - 0% 24,91 3,35 3230,04 288,76 71% 30% 17,21 1,02
11 917,18 - - - 0% 24,91 6,41 - - - 0% 18,35 4,23
The results presented in Table 13 suggest that the AFGR technique provides computational advantages, particularly in terms of model reduction and improved solver tractability. The introduction of the AFGR module leads to a systematic improvement in the average objective function value, with an average reduction in the optimality gap ranging between 2% and 10% across the tested instances. For instance, in the case with $V=6$, the gap decreases from 5% without AFGR to 2% with AFGR, highlighting the ability of AFGR to guide the search process toward higher-quality solutions. In terms of solution stability, the standard deviations of the objective values decrease significantly when AFGR is activated. This indicates greater robustness of the algorithm with respect to the inherent randomness of the heuristic optimization process. Although the addition of the AFGR mechanism introduces a slight increase in the average computational time, the total runtime remains substantially lower than that of exact approaches and is largely justified by the quality gains obtained. A more detailed analysis confirms that AFGR consistently reduces the average optimality gap across all problem sizes. Specifically, for $V=4$, the gap drops from 4% without AFGR to 1% with AFGR. For $V=6$, it decreases from 5% to 2%, and for $V=10$, a particularly challenging case, the improvement is even more pronounced, from 14% without AFGR to 6% with AFGR. Overall, the AFGR mechanism yields a gap reduction of approximately 40% to 70% compared to the baseline method, resulting in solutions that are significantly closer to the known optimum. Moreover, the percentage of successfully solved instances improves with the inclusion of AFGR, which confirms the greater efficiency of the adaptive repair mechanism in handling infeasible or borderline-feasible configurations. Integrating the AFGR module into the model without time windows leads to a substantial improvement in solution quality, enhanced robustness, and overall increased reliability of the heuristic procedure, while keeping the computational effort within acceptable limits.
The comparison be time window seen the models with and without time windows reveals several general trends regarding performance, solution quality, and computational behavior. The inclusion of time windows constraints significantly increases the problem complexity, as reflected by higher average objective function values and increased computational times, as we expected. On average, the presence of time windows leads to a deterioration of approximately 15–30% in solution quality, when comparing configurations with and without time windows under the same algorithmic settings. This performance drop is primarily due to the additional temporal feasibility constraints, which restrict the solution space and limit the flexibility of routing decisions. Despite the added difficulty introduced by time windows, the AFGR mechanism remains effective. Although the average optimality gap tends to be higher in the presence of time windows, as expected due to the stricter feasibility requirements and the reduced number of high-quality feasible solutions, AFGR consistently contributes to narrowing the gap with respect to the optimal benchmark (as computed by Gurobi). Moreover, the use of AFGR leads to a marked increase in the percentage of solved instances, particularly in the most constrained time windows cases. From a computational perspective, AFGR proves to be essential for recovering feasible solutions in time windows instances, where the baseline approach often either fails to find a solution or produces highly unstable outcomes. Execution times tend to increase further in time windows scenarios, especially in configurations without AFGR, which require more iterations and repair steps to reach feasibility. The AFGR enhanced configurations achieve a better trade-off be time window seen feasibility and runtime, supporting the robustness of the proposed mechanism in handling both temporal and spatial constraints simultaneously.
STSP-TW
This portion focuses on evaluating instances where the problem is limited to delivery tasks only, excluding pickup operations. The variant analyzed here corresponds to the configuration described in paragraph STSP-TWPD model without pickup operations in the subsection 6.4.1. We focus solely on the ABF formulation, rather than testing both ABF and NBF, since their equivalence has already been established in Section 4.3.1. The performance of the model is assessed under this configuration, considering both the presence and absence of the AFGR preprocessing strategy.
Table 14: Performance of the STSP-TW problem, evaluated with and without the application of AFGR. The table reports different instance size, defined by the number of vehicles $V$, and includes: OF opt: optimal OF value obtained with Gurobi, OF Avg: average OF value across ten runs, OF Std: standard deviation of the OF value, GAP: average optimality gap, calculated as the percentage deviation of OF Avg from OF opt, % Solved: percentage of instances solved successfully within the time limit, Time Avg: average solution time in seconds, Time Std: standard deviation of the solution time. A value of “–” indicates that no valid solution was obtained in any of the ten runs.
$V$ OFoptOF\>opt Without AFGR With AFGR
OF Avg OF Std GAP opt % Solved Time Avg Time Std OF Avg OF Std GAP opt % Solved Time Avg Time Std
4 882,8427 915,98 42,78 4% 100% 31,25 21,14 899,41 34,93 2% 100% 32,23 14,82
5 995,9593 1379,81 167,55 28% 100% 54,48 43,42 1286,42 195,08 23% 100% 39,15 20,65
6 1019,615 2077,03 130,95 51% 60% 30,48 20,75 2070,27 146,15 51% 80% 38,53 43,52
7 1028,379 2590,66 249,51 60% 30% 48,51 39,63 2572,72 206,69 60% 50% 45,62 33,47
8 1241,384 - - - 0% 35,69 12,13 2890,74 520,8 57% 40% 25,97 6,13
9 1299,211 - - - 0% 21,42 1,67 2471,66 71,11 47% 30% 19,99 1,86
10 1337,31 - - - 0% 36,38 6,08 - - - 0% 24,71 3,51
The results reported in Table 14 highlight the impact of excluding pickup operations on the behavior of the ABF model, both with and without the AFGR mechanism. For small-size instances ($V=4$), the performance is already strong even without AFGR: the average optimality gap is close to the reference value ($GAP=4\%$) with 100% of instances solved. Nevertheless, the inclusion of AFGR further improves the results by reducing the average gap to 2% and ensuring greater stability, as reflected in the lower standard deviation. In this case, the slight increase in average computational time is negligible and fully justified by the gain in robustness. When the problem size increases ($V=5$), the model becomes less stable: the gap rises to 28% without AFGR, with high variability across runs (Std=167,55). The introduction of AFGR reduces the average gap to 23% and increases solution consistency, although variability is not entirely eliminated. Moreover, the average computational time decreases significantly (from 54,48s to 39,15s). For $V=6$ and $V=7$, the problem becomes more challenging. Without AFGR, the gap reaches 51–60%, with only 60% and 30% of instances solved, respectively. The integration of AFGR does not reduce the gap, which remains high, but it considerably improves the percentage of solved instances (80% for $V=6$ and 50% for $V=7$) and leads to more stable outcomes, as shown by the reduction in standard deviation. This indicates that in more complex scenarios AFGR plays a crucial role primarily in ensuring feasibility and reducing variability rather than directly improving the objective function value. For larger instances ($V\geq 8$), the difficulty further increases. Without AFGR, the algorithm fails to produce any feasible solution (0% solved). With AFGR, non-optimal but consistent solutions are obtained: for $V=8$ the average gap is 57% with 40% of instances solved, while for $V=9$ the gap is 47% with 30% solved. These results confirm the essential role of AFGR as a recovery mechanism for feasibility in highly complex cases. Finally, for $V=10$, neither configuration is able to consistently generate valid solutions. The problem becomes intractable within the given time limits, suggesting that additional acceleration or decomposition strategies would be necessary to deal with instances of this scale.
Overall, the analysis shows that AFGR becomes increasingly beneficial as
7 Conclusions and Further Work
This work introduces the STSP-TWPD, a novel and practically relevant extension of the classical STSP. By integrating time windows, pickup and delivery operations and vehicle capacity constraints within a Steiner graph structure, the STSP-TWPD captures the multifaceted challenges faced in real-world logistics applications such as last-mile delivery, reverse logistics, and time-sensitive distribution. To address the inherent computational complexity of this NP-hard problem, we develop two original mathematical formulations: an ABF that extends traditional STSP models by explicitly incorporating temporal and logistical dimensions, and a NBF, a novel modeling We also introduce formal mathematical formulations for the STSP-PD and the STSP-TW, which, to the best of our knowledge, have not previously appeared in the literature. Furthermore, we propose a heuristic reduction method, the AFGR, a preprocessing technique that reduces problem dimensionality by eliminating redundant arcs, while preserving feasibility, significantly enhancing scalability and solver performance. We solved the proposed models to optimality using Gurobi and additionally assessed the feasibility and compatibility of the STSP-TWPD formulations within a hybrid quantum–classical computing framework. By implementing our models on D-Wave’s LeapCQMHybrid solver, we The experimental results confirm the effectiveness of the AFGR method in reducing model size by more than 40% on average, improving solution feasibility and stability across a wide range of configurations.
Looking ahead, several avenues for future research emerge from this work. From a modeling perspective, incorporating additional realistic constraints, such as stochastic travel times, multiple vehicles, dynamic request arrivals, or customer priority levels, would extend the applicability of the STSP-TWPD to more complex and operationally relevant logistics scenarios. A particularly meaningful direction is to relax the current assumption that each customer issues either a pickup or a delivery request, with pickups required to precede deliveries. Exploring more general pickup-and-delivery structures, such as simultaneous pickup and delivery, flexible precedence patterns, or multi-commodity flows, could significantly broaden the modeling capabilities of the STSP-TWPD framework.
On the computational side, this work also opens promising opportunities for advancing quantum-assisted optimization. Further investigation of hybrid quantum–classical solvers is warranted, including comparative analyses across quantum platforms, embedding techniques, and solver configurations. Such studies would help clarify how improvements in coherence time, qubit connectivity, and sampling strategies may enhance the scalability and performance of STSP-TWPD formulations in quantum settings.
Declaration of competing interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
Data availability
Data will be made available on request.
Acknowledgments
Eneko Osaba acknowledges support from the Basque Government through the ELKARTEK program, project "KUBIBIT - Kuantikaren Berrikuntzarako Ikasketa Teknologikoa" (KK-2025/00079).